### v2

# Scalable Feature Learning on Huge Knowledge Graphs for Downstream Machine Learning

Abstract
Many machine learning tasks can benefit from external knowledge. Large knowledge graphs store such knowledge, and embedding methods can be used to distill it into ready-to-use vector representations for downstream applications. For this purpose, current models have however two limitations: they are primarily optimized for link prediction, via local contrastive learning, and their application to the largest graphs requires significant engineering effort due to GPU memory limits. To address these, we introduce SEPAL: a Scalable Embedding Propagation ALgorithm for large knowledge graphs designed to produce high-quality embeddings for downstream tasks at scale. The key idea of SEPAL is to ensure global embedding consistency by optimizing embeddings only on a small core of entities, and then propagating them to the rest of the graph with message passing. We evaluate SEPAL on 7 large-scale knowledge graphs and 46 downstream machine learning tasks. Our results show that SEPAL significantly outperforms previous methods on downstream tasks. In addition, SEPAL scales up its base embedding model, enabling fitting huge knowledge graphs on commodity hardware. Our code is available at: https://github.com/flefebv/sepal.git.
1 Introduction: embedding knowledge for downstream tasks
External knowledge for machine learning
Bringing general knowledge to a machine-learning task revives an old promise of making it easier via this knowledge (lenat2000thresholds). Indeed, data science is often about entities of the world –persons, places, organizations– that are well characterized in general-purpose knowledge graphs. These graphs carry rich information, including numerical attributes and relationships between entities, and can be connected to string values in tabular data through entity linking techniques (mendes2011dbpedia; foppiano2020entity; delpeuch2019opentapioca). A thorny challenge, however, is to transform this relational information into features for downstream tabular machine learning (kanter2015deep; cappuzzo2024retrieve; robinson2025relbench). To that end, a scalable solution is offered by graph embedding methods that distill the graph information into node features readily usable by any downstream tabular learner (grover2016node2vec; cvetkov2023relational; ruiz2024high).
Knowledge graphs as general knowledge sources
The rapid growth of general-purpose knowledge graphs brings the exciting prospect of a very general feature enrichment. Indeed, a richer knowledge graph provides more comprehensive coverage and context, thereby bringing greater value to the downstream analysis (ruiz2024high). ConceptNet pioneered the distribution of general-knowledge embeddings, building on a graph of 8 million entities (speer2017conceptnet). Since then, knowledge graphs have continued to expand rapidly. For instance, as of 2025, Wikidata (vrandevcic2014wikidata) describes 115M entities and gains around 15M yearly (wikidatagrowth), and YAGO4 (pellissier2020yago) gives a curated view on 67M entities.
Optimizing embeddings for the right task
In parallel, the sophistication of knowledge-graph embedding (KGE) models is increasing (bordes2013translating; yang2015embedding; balazevic2019multi), capturing better the relational aspect of the data, important for downstream tasks (cvetkov2023relational). However, most of the KGE literature prioritizes link prediction as the primary benchmark, despite recent findings showing that strong performance on this task does not correlate with improved performance on downstream predictive tasks (ruffinelli2024beyond). One reason may be that, for link prediction, models typically optimize for local contrasts, resulting in embeddings that are not calibrated (Tabacof2020Probability; arakelyan2023adapting). While prior work has explored multi-hop reasoning to capture more complex graph patterns (hamilton2018embedding; ren2020beta), the standard evaluation paradigm for KGEs still revolves around internal tasks, rather than how the learned embeddings can transfer knowledge to practical machine-learning tasks beyond the knowledge graph itself.
The importance of scalability
To leverage the full potential of very large knowledge graphs, embedding methods must be highly scalable. While many methods have been proposed to scale KGE models, doing so is not trivial. Sophisticated KGE models are typically demonstrated on small datasets like FB15k (15k entities) or WN18 (40k entities), which are orders of magnitude smaller than modern general-purpose or industrial knowledge graphs (sullivan2020reintroduction). The common solution to this scalability challenge is either distributed computation across multiple GPUs or machines (lerer2019pytorch; zhu2019graphvite; zheng2020dgl; dong2022het; zheng2024ge2), or leveraging the full memory hierarchy (disk, CPU, and GPU) on a single machine (mohoney2021marius; ren2022smore). These approaches require significant engineering effort to manage data partitioning, optimize data movement, and minimize synchronization overheads.
Contributions
In this paper, we aim to bridge the gap between advances in embedding methods and the goal to create large and reusable general-knowledge embeddings for downstream applications. We introduce SEPAL, a scalable algorithm that applies as a wrapper to many embedding models. Our contributions are:
  1. 1.
    We propose a new embedding optimization strategy that enforces global consistency. Instead of optimizing all embeddings with local contrastive learning, SEPAL first processes a small but dense core of the graph, to learn relation and core-entity embeddings. It then propagates these embeddings to the remaining entities using relation-aware message passing. The absence of negative sampling at this propagation stage accelerates the embedding computation and makes them better suited for downstream tasks.
  2. 2.
    We provide a theoretical analysis showing that SEPAL’s propagation step, combined with DistMult, implicitly maximizes the alignment of embeddings within positive triples.
  3. 3.
    We introduce BLOCS, a scalable graph-splitting algorithm that partitions huge, scale-free graphs into manageable, overlapping subgraphs. This enables fitting the embedding process on a single GPU, avoiding the engineering complexity of distributed systems. Here, the challenge lies in the scale-free and connectivity properties of a large knowledge graph: some nodes are connected to a significant fraction of the graph, while others are hard to reach.
  4. 4.
    We conduct an extensive empirical study on 7 large knowledge graphs and 46 downstream tasks. Results show that SEPAL significantly outperforms standard methods on downstream tasks and is generally faster than existing large-scale systems. Moreover, it scales to ultra-large graphs with little computational resources: we embed WikiKG90Mv2 –91M entities, 601M triples– with a single 32GB V100 GPU. Our experimental results also highlight that using such large knowledge graphs is beneficial for downstream tasks in real-world feature enrichment scenarios.
We start by reviewing related work in section 2. Then, section 3 describes our contributed method and section 4 gives a theoretical analysis. In section 5 we evaluate SEPAL’s performance on knowledge graphs of increasing size between YAGO3 (2.6M entities, mahdisoltani2014yago3) and WikiKG90Mv2 (91M entities, hu2020open); we study the use of the embeddings for feature enrichment on 46 downstream machine learning tasks, showing that SEPAL makes embedding methods more tractable while generating better embeddings for downstream tasks. Finally, section 6 discusses the contributions and limitations of SEPAL.
2 Related work: embedding optimization and scalability
Knowledge graphs are multi-relational graphs storing information as triples $(h,r,t)$, where $h$ is the head entity, $r$ is the relation, and $t$ is the tail entity. We denote $\mathcal{V}$ and $\mathcal{R}$ respectively the set of entities and relations, and $\mathcal{K}$ the set of triples of a given knowledge graph ($\mathcal{K}\subset\mathcal{V}\times\mathcal{R}\times\mathcal{V}$).
2.1 Optimizing knowledge-graph embeddings
Here, we provide an overview of approaches to generate low-dimensional (typically $d=100$) vector representations for the entities, that can be used in downstream applications. These include both graph-embedding techniques, that leave aside the relations, and KGE methods accounting for relations.
Graph embedding
A first simple strategy to get very cost-effective vector representations is to compute random projections. This avoids relying on –potentially costly– optimization, and provides embeddings preserving the pairwise distances to within an epsilon (dasgupta2003elementary). FastRP (chen2019fast) proposes a scalable approach, with a few well-chosen very sparse random projections of the normalized adjacency matrix and its powers.
Another family of methods performs explicit matrix factorizations on matrices derived from the adjacency matrix, for instance GraREP (cao2015grarep) or NetMF (qiu2018network). These methods output close embedding vectors for nodes with similar neighborhoods.
Similarly, Skip-Gram Negative Sampling (SGNS), behind word2vec (mikolov2013efficient; mikolov2013distributed), performs an implicit factorization (levy2014neural). It has been adapted to graphs: DeepWalk (perozzi2014deepwalk) and node2vec (grover2016node2vec) use random walks on the graph to generate “sentences" fed to word2vec. LINE (tang2015line) explores a similar strategy varying edge sampling. Here, the loss function is typically a binary logistic regression objective:
$\mathcal{L}_{\text{SGNS}}=-\log\sigma(\boldsymbol{\theta}_{w_{c}}^{\top}\boldsymbol{\theta}_{w_{t}})-\sum_{i=1}^{p}\log\left(1-\sigma(\boldsymbol{\theta}_{w_{i}}^{\top}\boldsymbol{\theta}_{w_{t}})\right)$ (1)
with $\boldsymbol{\theta}_{w_{t}}$ and $\boldsymbol{\theta}_{w_{c}}$ the embeddings of the target and context nodes (or words), $w_{i}$ the $i$-th negative sample drawn from a noise distribution, $p$ the number of negative samples, and $\sigma$ the sigmoid function.
Table 1: Expression of $\phi$ in some embedding models. $\odot$ denotes the Hadamard product, $\otimes$ the Hamilton product, and $\times_{i}$ the tensor product along mode ii. The models we list here are all compatible with our proposed SEPAL approach.
Model Relational operator $\phi$
TransE (bordes2013translating) $\boldsymbol{\theta}_{h}+\boldsymbol{w}_{r}$
MuRE (balazevic2019multi) $\boldsymbol{\theta}_{h}\odot\boldsymbol{\rho}_{r}-\boldsymbol{w}_{r}$
RotatE (sun2019rotate) $\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$
QuatE (zhang2019quaternion) $\boldsymbol{\theta}_{h}\otimes\boldsymbol{w}_{r}$
DistMult (yang2015embedding) $\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$
ComplEx (trouillon2016complex) $\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$
TuckER (balavzevic2019tucker) $\boldsymbol{\mathcal{W}}\times_{1}\boldsymbol{\theta}_{h}\times_{2}\boldsymbol{w}_{r}$
Knowledge-graph embedding
RDF2vec (ristoski2016rdf2vec) adapts SGNS to multi-relational graphs by simply adding the relations to the generated sentences.
More advanced methods model relations as geometric transformations in the embedding space. These triple-based methods, inspired by SGNS, represent the plausibility of a triple given the embeddings $\boldsymbol{\theta}_{h},\boldsymbol{w}_{r},\boldsymbol{\theta}_{t}$ of the entities and relation with a scoring function $f(h,r,t)$ often written as
Scoring function $f(h,r,t)=-\mathrm{sim}(\phi(\boldsymbol{\theta}_{h},\boldsymbol{w}_{r}),\boldsymbol{\theta}_{t})$ (2)
where $\phi$ is a model-specific relational operator, and $\mathrm{sim}$ a similarity function. The embeddings are optimized by gradient descent to maximize the score of positive triples, and minimize that of negative ones. A possible loss function is the binary cross-entropy loss (ali2021bringing)
$\mathcal{L}_{\text{BCE}}=-\log\sigma(f(h,r,t))-\sum_{i=1}^{p}\log\big(1-\sigma(f(h_{i}^{\prime},r,t_{i}^{\prime}))\big)$ (3)
which boils down to SGNS for $f(h,r,t)=\boldsymbol{\theta}_{h}^{\top}\boldsymbol{\theta}_{t}$. These models strive to align, for positive triples, the tail embedding $\boldsymbol{\theta}_{t}$ with the “relationally” transformed head embedding $\phi(\boldsymbol{\theta}_{h},\boldsymbol{w}_{r})$. The challenge is to design a clever $\phi$ operator to model complex patterns in the data, like hierarchies, compositions, or symmetries. Indeed some relations are one-to-one (people only have one biological mother), well represented by a translation (bordes2013translating), while others are many-to-one (for instance many person were BornIn Paris), calling for $\phi$ to contract distances (wang2017knowledge). Many models explore different parametrizations, among which MuRE (balazevic2019multi), RotatE (sun2019rotate), or QuatE (zhang2019quaternion) have good performance (ali2021bringing). This framework also includes models like DistMult (yang2015embedding), ComplEx (trouillon2016complex), or TuckER (balavzevic2019tucker), that implicitly perform tensor factorizations.
Embedding propagation
To smooth computed embeddings, CompGCN (vashishth2020compositionbased) introduces the idea of propagating knowledge-graph embeddings using the relational operator $\phi$, but couples it with learnable weights and a non-linearity. REP (wang2022rep) simplifies this framework by removing weight matrices and non-linearities. rossi2022unreasonable also use Feature Propagation, but to impute missing node features in graphs. albooyeh2020out incorporate propagation within the standard link prediction pipeline, with negative sampling and gradient descent on standard KGE loss functions.
2.2 Techniques for scaling graph algorithms
Various tricks help scale graph algorithms to the sizes we are interested in –millions of nodes.
Graph partitioning
Scaling up graph computations, for graph embedding or more generally, often relies on breaking down graphs in subgraphs. Section˜E.1 presents corresponding prior work.
Local subsampling
Other forms of data reduction can help to scale graph algorithms (e.g. based on message passing). Algorithms may subsample neighborhoods, as GraphSAGE (hamilton2017inductive) that selects a fixed number of neighbors for each node on each layer, or MariusGNN (waleffe2023mariusgnn) that uses an optimized data structure for neighbor sampling and GNN aggregation. Cluster-GCN (chiang2019cluster) restricts the neighborhood search within clusters, obtained by classic clustering algorithms, to improve computational efficiency. GraphSAINT (zeng2020graphsaint) samples overlapping subgraphs through random walks, for supervised GNN training via node classification, optimizing GNN weights without processing the full graph.
Multi-level techniques
Multi-level approaches, such as HARP (chen2018harp), GraphZoom (deng2019graphzoom) or MILE (liang2021mile), coarsen the graph, compute embeddings on the smaller graph, and project them back to the original graph.
2.3 Scaling knowledge-graph embedding
Parallel training
Many approaches scale triple-level stochastic solvers by distributing training across multiple workers, starting from the seminal PyTorch-BigGraph (PBG) (lerer2019pytorch) that splits the triples into buckets based on the partitioning of the entities. The challenge is then to limit overheads and communication costs coming from 1) the additional data movement incurred by embeddings of entities occurring in several buckets 2) the synchronization of global trainable parameters such as the relation embeddings. For this, DGL-KE (zheng2020dgl) reduces data movement by using sparse relation embeddings and METIS graph partitioning (karypis1997metis) to distribute the triples across workers. HET-KG (dong2022het) further optimizes distributed training by preserving a copy of the most frequently used embeddings on each worker, to reduce communication costs. These “hot-embeddings" are periodically synchronized to minimize inconsistency. SMORE (ren2022smore) leverages asynchronous scheduling to overlap CPU-based data sampling, with GPU-based embedding computations. Algorithmically, it contributes a rejection sampling strategy to generate the negatives at low cost. GraphVite (zhu2019graphvite) accelerates SGNS for graph embedding by both parallelizing random walk sampling on multiple CPUs, and negative sampling on multiple GPUs. Marius (mohoney2021marius) reduces synchronization overheads by opting for asynchronous training of entity embeddings with bounded staleness, and minimizes IO with partition caching and buffer-aware data ordering. GE2 (zheng2024ge2) improves data swap between CPU and multiple GPUs. Finally, the LibKGE (libkge) Python library also supports parallel training and includes GraSH (kochsiek2022start), an efficient hyperparameter optimization algorithm for large-scale KGE models.
Parameter-efficient methods
Other approaches reduce GPU memory pressure by limiting the number of parameters. NodePiece (galkin2021nodepiece) and EARL (chen2023entity) embed a subset of entities and train an encoder to compute the embeddings of the other entities. However, their tokenization step is costly, and they have not been demonstrated on graphs larger than 2.5M nodes.
3 SEPAL: revisiting knowledge-graph embedding optimization
Figure 1: SEPAL’s embedding pipeline. First, a core subgraph is extracted from the input knowledge graph (step 1.a). BLOCS then subdivides this input knowledge graph into outer subgraphs (step 1.b). Next, the core subgraph is embedded using traditional KGE models, which generate vector representations for both core entities and relations (step 2.). Finally, these embeddings are propagated with message passing to each outer subgraph successively (step 3.).
Most work on scaling knowledge-graph embedding has focused on efficient parallel computing to speed up stochastic optimization. We introduce a different approach, SEPAL, which changes how embeddings are computed, enforcing a more global structure, beneficial for downstream tasks, while avoiding much of the optimization cost. To that end, SEPAL proceeds in three steps (Figure 1):
  1. 1.
    separates the graph in a core and a set of connected overlapping outer subgraphs that cover the full graph;
  2. 2.
    uses a classic KGE model to optimize the embeddings of the core entities and relations;
  3. 3.
    propagates the embeddings from the core to the outer subgraphs, using a message-passing strategy preserving the relational geometry and ensuring global embedding consistency, with no further training.
SEPAL’s key idea is to allocate more computation time to the more frequent entities and then use message passing to propagate embeddings at low cost to regions of the graph where they have not been computed yet. It departs from existing embedding propagation methods (vashishth2020compositionbased; wang2022rep) that compute embeddings on the full graph and use propagation as post-processing to smooth them. SEPAL is compatible with any embedding model whose scoring function has the form given by Equation 2, examples of which are provided in Table 1.
3.1 Splitting large graphs into manageable subgraphs
Breaking up the graph into subgraphs is key to scaling up our approach memory-wise. Specifically, we seek a set of subgraphs that altogether cover the full graph but are individually small enough to fit on GPUs, to enable the subsequent GPU-based message passing.
Core subgraph
SEPAL first defines the core of a knowledge graph. The quality of the core embeddings is particularly important, as they serve as (fixed) boundary conditions during the propagation phase. Good relation embeddings are also key to structuring the propagation. To optimize this quality, two key factors must be considered during core selection: 1) ensuring a dense core subgraph by selecting the most central entities and 2) achieving full coverage of relation types. Yet, there can be a trade-off between these two objectives, hence, SEPAL offers two core selection procedures:
Degree-based selection: This simple approach selects the top entities by degree –with proportion $\eta_{n}\in(0,1)$– and keeps only the largest connected component of the induced subgraph. The resulting core is dense, which boosts performance for entity-centric tasks like feature enrichment (Section˜F.2.2). However, it does not necessarily contain all the relation types.
Hybrid selection: To ensure full relational coverage, this method combines two sampling strategies. First, it selects entities with the highest $\eta_{n}$ degrees. Second, for each relation type, it includes entities involved in edges with the highest $\eta_{e}$ degrees (where degree is the sum of the head and tail nodes’ degrees). The union of these two sets forms the core, and if disconnected, SEPAL reconnects it by adding the necessary entities (details in Section˜F.2.2). Hyperparameters $\eta_{n},\eta_{e}\in(0,1)$ are proportions of nodes and edges that control the core size.
Compared to degree-based selection, hybrid selection benefits tasks relying on relation embeddings, such as link prediction. However, its additional relation-specific edge sampling and reconnection steps can be computationally expensive for knowledge graphs with many relations. For disconnected input graphs, all connected components other than the largest one are added to the core subgraph.
Outer subgraphs
The next class of subgraphs that we generate –the outer subgraphs– aim at covering the rest of the graph. The purpose of these subgraphs demands the following requirements:
R1: connected
the subgraphs must be connected, to propagate the embeddings;
R2: bounded size
the subgraphs must have bounded sizes, to fit their embeddings in GPU memory;
R3: coverage
the union of the subgraphs must be the full graph, to embed every entity;
R4: scalability
extraction must run with available computing resources, in particular memory.
Extracting such subgraphs is challenging on large knowledge graphs. These are scale-free graphs with millions of nodes and no well-defined clusters (leskovec2009community). They pose difficulties to partitioning algorithms. For instance, algorithms based on propagation, eigenvalues, or power iterations of the adjacency matrix (raghavan2007near; shi2000normalized; newman2006finding) struggle with the presence of extremely high-degree nodes that make the adjacency matrix ill-conditioned. None of the existing partitioning algorithms satisfy our full set of constraints, and thus we devise our own algorithm, called BLOCS and described in detail in Appendix˜E. To satisfy the requirements despite these challenges, BLOCS creates overlapping subgraphs.
We contribute BLOCS, an algorithm designed to break large graphs into Balanced Local Overlapping Connected Subgraphs. The name summarizes the goals: 1) Balanced: BLOCS produces subgraphs of comparable sizes. $m$, the upper bound for subgraph sizes, is a hyperparameter. 2) Local: the subgraphs have small diameters. This locality property is important for the efficiency of SEPAL’s propagation phase, as it reduces the number of propagation iterations needed to converge to the global embedding structure. 3) Overlapping: a given node can belong to several subgraphs. This serves our purpose because it facilitates information transfer between the different subgraphs during the propagation. 4) Connected: all generated subgraphs are connected.
BLOCS uses three base mechanisms to grow the subgraphs: diffuse (add all neighboring entities to the current subgraph), merge (merge two overlapping subgraphs), and dilate (add all unassigned neighboring entities to the current subgraph). There are two different regimes during the generation of subgraphs. First, few entities are assigned, and the computationally effective diffusion quickly covers a large part of the graph, especially entities that are close to high-degree nodes. However, once these close entities have been assigned, the effectiveness of diffusion drops because it struggles to reach entities farther away. For this reason, BLOCS switches from diffusion to dilation once the proportion of assigned entities reaches a certain threshold $h$ (a hyperparameter chosen $\approx.6$, depending on the dataset). By adding only unassigned neighbors to subgraphs, dilation drives subgraph growth towards unassigned distant entities. However, the presence of long chains can drastically slow down this regime because they make it add entities one by one. Some knowledge graphs have long chains, for instance YAGO4.5 (see Diameter in Table 2). To tackle them, BLOCS switches back to diffusion for a few steps, with seeds taken inside the long chains.
BLOCS works faster on graphs that have small diameters, where most entities can be reached during the diffusion regime and fewer dilation steps are required (Section˜F.1).
3.2 Core optimization with traditional KGE models
Once the core subgraph is defined, SEPAL trains on GPU any compatible triple-based embedding model (DistMult, TransE, …). This process generates embeddings for the core entities and relations, including inverse relations, added to ensure connectedness for the subsequent propagation step.
3.3 Outside the core: relation-aware propagation
Key to SEPAL’s global consistency of embeddings and to computational efficiency is that it does not use contrastive learning and gradient descent for the outer entities. Instead, the final step involves an embedding propagation that is consistent with the KGE model (multiplication for DistMult, addition for TransE, …) and preserves the relational geometry of the embedding space. To do so, SEPAL leverages the entity-relation composition function $\phi$ (given by Table 1) used by the KGE model, and the embeddings of the relations $\boldsymbol{w}_{r}$ trained on the core subgraph. From Equation 2 one can derive, for a given triple $(h,r,t)$, the closed-form expression of the tail embedding that maximizes the scoring function $\operatorname*{arg\,max}_{\boldsymbol{\theta}_{t}}f(h,r,t)=\phi(\boldsymbol{\theta}_{h},\boldsymbol{w}_{r})$. SEPAL uses this property to compute outer embeddings as consistent with the core as possible, by propagating from core entities with message passing.
First, the embeddings are initialized with $\boldsymbol{\theta}_{u}^{(0)}=\begin{cases}\boldsymbol{\theta}_{u},&\text{if entity $u$ belongs the core subgraph,}\\ \textbf{0},&\text{otherwise.}\end{cases}$
Then, each outer subgraph $\mathcal{S}\subset\mathcal{K}$ is loaded on GPU, merged with the core subgraph $\mathcal{C}$, and SEPAL performs $T$ steps of propagation ($T$ is a hyperparameter), with the following message-passing equations:
Message: $\boldsymbol{m}_{v,u}^{(t+1)}$ $=\sum_{(v,r,u)\in\mathcal{S}\cup\mathcal{C}}\phi(\boldsymbol{\theta}_{v}^{(t)},\boldsymbol{w}_{r})\qquad$ (4)
Aggregation: $\boldsymbol{a}_{u}^{(t+1)}$ $=\sum_{v\in\mathcal{N}(u)}\boldsymbol{m}_{v,u}^{(t+1)}\qquad$ (5)
Update: $\boldsymbol{\theta}_{u}^{(t+1)}$ $=\frac{\boldsymbol{\theta}_{u}^{(t)}+\alpha\boldsymbol{a}_{u}^{(t+1)}}{\left\|\boldsymbol{\theta}_{u}^{(t)}+\alpha\boldsymbol{a}_{u}^{(t+1)}\right\|_{2}}$ (6)
where $\mathcal{N}(u)$ denotes the set of neighbors of outer entity $u$, $\mathcal{K}$ the set of positive triples of the graph, and $\alpha$ a hyperparameter similar to a learning rate. During updates, $\ell_{2}$ normalization projects embeddings on the unit sphere. With DistMult, this accelerates convergence by canceling the effect of neighbors that still have zero embeddings. Normalizing embeddings is a common practice in knowledge-graph embedding (bordes2013translating; yang2015embedding), and SEPAL acts accordingly. During propagation, the core embeddings remain frozen.
4 Theoretical analysis: embedding alignment
SEPAL minimizes a global energy via gradient descent
Proposition˜4.1 shows that SEPAL with DistMult minimizes an energy that only accounts for the positive triples. The more aligned the embeddings within positive triples, the lower this energy. In self-supervised learning, negative sampling is needed to prevent embeddings from collapsing to a single point (hafidi2022negative). However, in our case, this oversmoothing is avoided thanks to the boundary conditions of fixed core-entities and relations embeddings, which act as “anchors” in the embedding space.
Proposition 4.1 (Implicit Gradient Descent).
Let $\mathcal{E}$ be the “alignment energy” defined as
$\mathcal{E}=-\sum_{(h,r,t)\in\mathcal{K}}\left\langle\boldsymbol{\theta}_{t},\phi(\boldsymbol{\theta}_{h},\boldsymbol{w}_{r})\right\rangle,$ (7)
with $\phi(\boldsymbol{\theta}_{h},\boldsymbol{w}_{r})=\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$ being the DistMult relational operator. Then, SEPAL’s propagation step amounts to a mini-batch projected gradient step descending $\mathcal{E}$ under the following conditions:
  1. 1.
    SEPAL uses DistMult as base KGE model;
  2. 2.
    the embeddings of relations and core entities remain fixed, and serve as boundary conditions;
  3. 3.
    the outer subgraphs act as mini-batches.
As a consequence, SEPAL converges towards a stationary point of $\mathcal{E}$ on the unit sphere.
Proof sketch.
For an outer entity $u$, the gradient update of its embedding $\boldsymbol{\theta}_{u}^{(t+1)}=\boldsymbol{\theta}_{u}^{(t)}-\eta\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}_{u}^{(t)}}$ boils down to SEPAL’s propagation equation, for a learning rate $\eta=\alpha$. See Section˜D.2 for detailed derivations. ∎
Analogy to eigenvalue problems
Section˜D.1 shows that SEPAL’s propagation, when combined with DistMult, resembles an Arnoldi-like iteration, suggesting that entity embeddings converge toward generalized eigenvectors of an operator that integrates both global graph structure and relation semantics (via fixed relation embeddings).
Queriability of embeddings
Assume that each entity $u\in\mathcal{V}$ is associated with a set of scalar features $\boldsymbol{x}_{u}\in\mathbb{R}^{m}$, such that $\boldsymbol{x}_{u,i}$ denotes the $i$-th property of $u$. We say that the embeddings support queriability with respect to property $i\in{1,\dots,m}$ if,
$\exists f_{i}\in\mathcal{F}\quad\text{such that}\quad f_{i}(\boldsymbol{\theta}_{u})\approx\boldsymbol{x}_{u,i},$ (8)
where $\mathcal{F}$ denotes a class of functions from $\mathbb{R}^{d}$ to $\mathbb{R}$ constrained by some regularity conditions.This queriability property is key to the embeddings’ utility to downstream machine learning tasks.
In knowledge graphs, properties $\boldsymbol{x}_{u,i}$ that are explicitly represented by triples $(u,r_{i},v_{i})$ can be recovered via the scoring function $\phi$ used in the model, under the condition of well-aligned embeddings. Specifically, when such a relation $r_{i}$ exists, a natural querying function is
$f_{i}:\boldsymbol{x}\mapsto\phi(\boldsymbol{x},\boldsymbol{w}_{r_{i}})$ (9)
where $\boldsymbol{w}_{r_{i}}$ is the embedding of relation $r_{i}$ (we assume scalar embeddings for simplicity).
Proposition˜4.1 shows that SEPAL minimizes an energy that promotes global alignment between embeddings of positive triples, which we argue leads to embeddings with high queriability. In contrast, classic KGE methods use negative sampling to incorporate a supplementary constraint of local contrast between positive and negative triples. This adds discriminative power to the model, useful for link prediction, but can be detrimental to the global alignment of embeddings, which is important for queriability.
5 Experimental study: utility to downstream tasks
Knowledge graph datasets
To compare large knowledge graphs of different sizes, we use Freebase (bollacker2008freebase), WikiKG90Mv2 ((an extract of Wikidata) hu2020open), and three generations of YAGO: YAGO3 (mahdisoltani2014yago3), YAGO4 (pellissier2020yago), and YAGO4.5 (suchanek2024yago). We expand YAGO4 and YAGO4.5 into a larger version containing also the taxonomy, i.e., types and classes –which algorithms will treat as entities– and their relations. We discard numerical attributes and keep only the largest connected component (Section˜A.1). To perform an ablation study of SEPAL without BLOCS for which we need smaller datasets, we also introduce Mini YAGO3, a subset of YAGO3 built with the 5% most frequent entities. Knowledge graph sizes are reported in Figure 5.
On real-world downstream regression tasks
We evaluate the embeddings as node features in downstream tasks (grover2016node2vec; cvetkov2023relational; robinson2025relbench; ruiz2024high). Specifically, we incorporate the embeddings in tables as extra features and measure the prediction improvement of a standard estimator in regression tasks. This setup allows us to compare the utility of knowledge graphs of varying sizes. Indeed, for a task, a suboptimal embedding of a larger knowledge graph may be more interesting than a high-quality embedding of a smaller knowledge graph because the larger graph brings richer information, on more entities. We benchmark 4 downstream regression tasks (adapted from cvetkov2023relational): Movie revenues, US accidents, US elections, and housing prices. Details are provided in Section˜A.2.
Figure 2: Statistical performance on real-world tables. a) Pareto frontiers of averaged normalized prediction scores with respect to embedding times (log-scale). b) Critical difference diagrams (terpilowski2019scikit) of average ranks among the three methods (SEPAL, PBG and DGL-KE) that scale to every knowledge-graph dataset. The ranks are averaged over all tasks; a task being defined as the combination of a downstream table and a source knowledge graph. SEPAL gets the best average downstream performance for each of the 7 source knowledge graphs. Figure 5 gives the detailed results for each table. Appendix B.1 details the metric used.
Larger knowledge graphs do bring value (Figure 2), as they cover more entities of the downstream tasks (Figure 4). For each source graph, SEPAL provides the best embeddings and is much more scalable (details in Figure 5). Considering performance/cost Pareto optimality across methods and source graphs (Figure 2a), SEPAL achieves the best performance with reduced cost, but the simple baseline FastRP also gives Pareto-optimal results, for smaller costs. Although FastRP discards the type of relation, it performs better than most dedicated KGE methods. Its iterations also solve a more global problem, like SEPAL (Section˜D.1).
We used DistMult as base model as it is a classic and good performer (Ruffinelli2020You; Kadlec2017KnowledgeBC; jain2020knowledge). Figure 22 shows that SEPAL can speed up DistMult over 20 times for a given training configuration. For other scoring functions like RotatE and TransE, Figure 7 shows that SEPAL also improves the downstream performance of its base model.
On WikiDBs tables
We also evaluate SEPAL on tables from WikiDBs (vogel2024wikidbs), a corpus of databases extracted from Wikidata. We build 42 downstream tables (26 classification tasks, and 16 regression tasks), described in Section˜A.2. Four of them are used as validation tasks, to tune hyperparameters, and the remaining 38 are used as test tables. Figure 3 presents the aggregated results on these 38 test tables, showing that, here also, applying SEPAL to very large graphs provides the best embeddings for downstream node regression and classification, and that SEPAL brings a decreased computational cost. Regarding the performance-cost tradeoff, FastRP is once again Pareto-optimal for small computation times, highlighting the benefits of global methods.
Figure 3: Statistical performance on WikiDBs tables. a) Critical difference diagrams of scalable methods. Black lines connect methods that are not significantly different. b) Pareto frontiers of averaged normalized prediction scores with respect to embedding times (log-scale). Figure 6 in Section˜C.2 provides the detailed results for each of the 38 test tables.
6 Discussion and conclusion
Benefits of larger graphs
We have studied how to build general-knowledge feature enrichment from huge knowledge graphs. For this purpose, we have introduced a scalable method that captures more of the global structure than the classic KGE methods. Our results show the benefit of embedding larger graphs. There are two reasons to this benefit: (1) larger knowledge graphs can result in larger coverage of downstream entities (another important factor for this is the age of the dataset: recent ones have better coverage) (2) for two knowledge graphs with equal coverage, the larger one can result in richer representations because the covered entities have more context to learn from.
Limitations
Our work focuses on embeddings for feature enrichment of downstream tables, an active research field (cvetkov2023relational; ruiz2024high; robinson2025relbench). Another popular use case for embeddings is knowledge graph completion. However, this task is fundamentally different from feature enrichment: embeddings optimized for link prediction may not perform well for feature enrichment, and vice versa (ruffinelli2024beyond). Nevertheless, we also evaluate SEPAL on knowledge graph completion, for which we expect lower performances given that SEPAL does not locally optimize the contrast between positive and negative triples scores. Results reported in Section˜C.6 show that SEPAL sometimes performs lower than existing methods (DGL-KE, PBG) on this task, although no method is consistently better than the others for all datasets, and statistical tests show no significant differences (Figure 8).
Conclusion: embeddings for downstream machine learning
In this paper, we show how to optimize knowledge-graph embeddings for downstream machine learning. We propose a highly scalable method, SEPAL, and conduct a comprehensive evaluation on 7 knowledge graphs and 46 downstream tables showing that SEPAL both: (1) markedly improves predictive performance on downstream tasks and (2) brings computational-performance benefits –multiple-fold decreased train times and bounded memory usage– when embedding large-scale knowledge graphs. Our theoretical analysis suggests that SEPAL’s strong performance on downstream tasks stems from its global optimization approach, resulting in better-aligned embeddings compared to classic methods based on negative sampling. SEPAL improves the quality of the generated node features when used for data enrichment in external (downstream) tasks, a setting that can strongly benefit from pre-training embeddings on knowledge bases as large as possible. It achieves this without requiring heavy engineering, such as distributed computing, and can easily be adapted to most KGE models.
Insights brought by our experiments go further than SEPAL. First, the method successfully exploits the asymmetry of information between “central” entities and more peripheral ones. Power-law distributions are indeed present on many types of objects, from words (piantadosi2014zipf) to geographical entities (giesen2011zipf) and should probably be exploited for general-knowledge representations such as knowledge-graph embeddings. Second, and related, breaking up large knowledge graphs in communities is surprisingly difficult: some entities just belong in many (all?) communities, and others are really hard to reach. Our BLOCS algorithm can be useful for other graph engineering tasks, such as scaling message-passing algorithms or simply generating partitions. Finally, the embedding propagation in SEPAL appears powerful, and we conjecture it will benefit further approaches. First, it can be combined with much of the prior art to scale knowledge-based embedding. Second, it could naturally adapt to continual learning settings (vandeven2022three; hadsell2020embracing; biswas2023knowledge), which are important in knowledge-graph applications since knowledge graphs, such as Wikidata, are often continuously updated with new information (Section˜G.3).
Acknowledgements
GV acknowledges support from ANR via grant TaFoMo (ANR-25-CE23-1822). This work is partly supported by Hi! PARIS and ANR/France 2030 program (ANR-23-IACL-0005).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope. Both abstract and introduction clearly articulate the paper’s contributions.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Section˜6 discusses limitations.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Section˜D.2 provides a complete proof of the main theoretical result, and discusses the full set of assumptions.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All the code necessary to reproduce the results is provided in an anonymized zip file and will be released by the time of publication.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All the code is provided in an anonymized zip file and will be released by the time of publication.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The experimental setting is presented in detail in appendix.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Critical difference diagrams associated with statistical significance tests are provided with the results.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: The experimental setup is fully described in appendix.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research conducted in the paper conforms with the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Section˜G.4 discusses potential societal impacts of the work performed.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All original papers are cited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: The paper does not release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: The core method development in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
Appendix A Datasets
A.1 Statistics on knowledge graph datasets
Table 2: Additional statistics on the knowledge graph datasets used. MSPL stands for Mean Shortest Path Length. The LCC column gives the percentage of entities of the graph that are in the largest connected component.
Maximum degree Average degree MSPL Diameter Density LCC
Mini YAGO3 65 711 12.6 3.3 11 1e-4 99.98%
YAGO3 934 599 4.0 4.2 23 2e-6 97.6%
YAGO4.5 6 434 121 4.5 5.0 502 1e-7 99.7%
YAGO4.5+T 6 434 122 5.0 4.0 5 1e-7 100%
YAGO4 8 606 980 12.9 4.5 28 3e-7 99.0%
YAGO4+T 32 127 569 9.4 3.4 6 1e-7 100%
Freebase 10 754 238 4.9 4.7 100 6e-8 99.1%
WikiKG90Mv2 37 254 176 12.8 3.6 98 1e-7 100.0%
More statistics on the knowledge graph datasets are given in Table 2. Maximum and average degree figures highlight the scale-free nature of real-world knowledge graphs. The values for mean shortest path length (MSPL) and diameter (the diameter is the longest shortest path) are provided for the largest connected component (LCC). They are remarkably small, given the number of entities in the graphs. Contrary to other datasets, YAGO4.5, Freebase, and WikiKG90Mv2 contain ‘long chains’ of nodes, which account for their larger diameters.
The density $D$ is the ratio between the number of edges $|E|$ and the maximum possible number of edges:
$D=\frac{|E|}{|\mathcal{V}|(|\mathcal{V}|-1)}$
where $|\mathcal{V}|$ denotes the number of nodes.
The LCC statistics show that for each knowledge graph, the largest connected component regroups almost all the entities.
A.2 Downstream tables
Real-world tables
We use 4 real-world downstream tasks adapted from cvetkov2023relational who also investigate knowledge-graph embeddings to facilitate machine learning. The specific target values predicted for each dataset are the following:
US elections:
predict the number of votes per party in US counties;
Housing prices:
predict the average housing price in US cities;
US accidents:
predict the number of accidents in US cities;
Movie revenues:
predict the box-office revenues of movies.
For each table, a log transformation is applied to the target values as a preprocessing step. Table 3 contains the sizes of these real-world downstream tables.
Table 3: Number of rows in the downstream tables.
US elections Housing prices US accidents Movie revenues
Number of rows 13 656 22 250 20 332 7 398
WikiDBs tables
WikiDBs contains 100,000 databases (collections of related tables), which altogether include 1,610,907 tables. We extracted 42 of those tables to evaluate embeddings. Here we describe the procedure used for table selection and processing.
Table selection: Most of the WikiDBs tables are very small –typically a few dozen samples– so our first filtering criterion was the table size, which must be large enough to enable fitting an estimator. Therefore, we randomly sampled 100 tables from WikiDBs with sufficient sizes ($N_{\text{rows}}>1,000$). Then we looked at each sampled table individually and kept those that could be used to define a relevant machine learning task (either regression or classification). We ended up with 16 regression tasks and 26 classification tasks.
Table processing: We removed rows with missing values, and reduced the size of large tables to keep evaluation tractable. For regression tables, we simply sampled 3,000 rows randomly (if the table had more than 3,000). We applied a log transformation to the target values to remove the skewness of their distributions. Before that, dates were converted into floats (by first converting them to fractional years, and then applying the transform $t\mapsto 2025-t$). For classification tables, to reduce the dataset sizes while preserving both class diversity and balance, we downsampled the tables with the following procedure:
  1. 1.
    Class filtering: we set a threshold $r=\min(50,0.9N_{2})$, where $N_{2}$ is the cardinality of the second most populated class, and retained only the classes with more than $r$ occurrences.
  2. 2.
    Limit number of classes: if more than 30 classes remained after filtering, only the 30 most frequent were kept.
  3. 3.
    Downsampling: if the resulting table contained more than 3,000 rows, we sampled rows such that: (a) if $r\cdot N_{\text{classes}}\leq 3,000$, at least $r$ rows were sampled per class, (b) if $r\cdot N_{\text{classes}}>3,000$, we sampled an approximately equal number of rows per class, fitting within the 3,000-row limit.
The specifications of the 42 tables extracted are given in Table 4 and Table 5.
Table 4: Regression tables from WikiDBs. The ‘DB number’ is the number of the database in WikiDBs from which the table was taken. Among these 16 regression tables, 2 are used for validation, and 14 for test.
Table name DB number Value to predict $\textbf{N}_{\text{rows}}$ Set
Historical Figures 62 826 Birth date 3 000 Val
Geopolitical Regions 66 610 Land area 2 324 Val
Eclipsing Binary Star Instances 3 977 Apparent magnitude 3 000 Test
Research Article Citations 14 012 Publication date 3 000 Test
Drawings Catalog 14 976 Artwork height 3 000 Test
Municipal District Capitals 19 664 Population count 2 846 Test
Twinned Cities 28 146 Population 1 194 Test
Ukrainian Village Instances 28 324 Elevation (meters) 3 000 Test
Dissolved Municipality Records 46 159 Dissolution date 3 000 Test
Research Articles 53 353 Publication date 3 000 Test
Territorial Entities 82 939 Population count 3 000 Test
Artworks Inventory 88 197 Artwork width 3 000 Test
Business Entity Locations 89 039 Population count 3 000 Test
WWI Personnel Profiles 89 439 Birth date 3 000 Test
Registered Ships 90 930 Gross tonnage 3 000 Test
Poet Profiles 94 062 Death date 3 000 Test
Table 5: Classification tables from WikiDBs. The ‘DB number’ is the number of the database in WikiDBs from which the table was taken. Among these 26 classification tables, 2 are used for validation, and 24 for test.
Table name DB number Class to predict $\textbf{N}_{\text{rows}}$ $\textbf{N}_{\text{classes}}$ Set
Creative Commons Authors 9 510 Gender 2 999 2 Val
Historical Figures 73 376 Profession 1 044 5 Val
Historic Buildings 473 Country 2 985 30 Test
Striatum Scientific Articles 2 053 Journal name 2 986 30 Test
Researcher Profile 7 136 Affiliated institution 237 7 Test
Decommissioned Transport Stations 7 310 Country 2 983 30 Test
Artist Copyright Representation 7 900 Artist occupation 2 986 27 Test
Forward Players 15 542 Team 2 985 30 Test
Rafael Individuals 29 832 Nationality 2 966 12 Test
Artworks Catalog 30 417 Artwork type 2 991 17 Test
Magic Narrative Motifs 36 100 Cultural origin 2 993 12 Test
Geographer Profiles 42 562 Language 2 992 14 Test
Surname Details 47 746 Language 1 420 5 Test
Sculpture Instances 56 474 Material used 2 985 30 Test
Spring Locations 63 797 Country 2 981 30 Test
Noble Individuals 64 477 Role 2 987 30 Test
Defender Profiles 65 102 Defender position 2 998 5 Test
Kindergarten Locations 66 643 Country 2 998 4 Test
Sub Post Office Details 67 195 Administrative territory 2 986 30 Test
State School Details 70 780 Country 2 995 12 Test
Notable Trees Information 70 942 Tree species 2 992 19 Test
Parish Church Details 87 283 Country 2 993 15 Test
Museum Details 90 741 Country 2 986 30 Test
Island Details 92 415 Country 2 986 30 Test
Philosopher Profiles 97 229 Language 2 985 29 Test
Music Albums Published in the US 97 297 Music Genre 2 984 30 Test
A.3 Entity coverage of downstream tables
Figure 4: Entity coverage of downstream tables. Over the 46 downstream tables 4 are used for validation (in blue), and 42 are used for test (in maroon).
Entity coverage
We define the coverage of table $T$ by knowledge graph $K$ as the proportion of downstream entities in $T$ that are described in $K$. Figure 4 gives the empirically measured coverage of the 46 downstream tables by the 8 different knowledge graphs used in our experimental study. It shows that larger and more recent knowledge graphs yield greater coverage.
Entity matching
Leveraging knowledge-graph embeddings to enrich a downstream tabular prediction task requires mapping the table entries to entities of the knowledge graph. We call this process entity matching.
For the 4 real-world tables, we performed the entity matching ‘semi-automatically’, following cvetkov2023relational. The entries of these tables are well formatted and a small set of simple rules is sufficient to match the vast majority of entities. Human supervision was required for some cases of homonymy, for instance.
For the WikiDBs tables, we used the Wikidata Q identifiers (QIDs) included in the WikiDBs dataset to straightforwardly match the entities to WikiKG90Mv2111Entity mapping for WikiKG90Mv2 is provided here., YAGO4, and YAGO4.5, which all provide the QIDs for every entity. YAGO4 also offers a mapping to the Freebase entities, which we used to match the Freebase entities to the WikiDBs tables. Additionally, both YAGO3 and YAGO4 provide mappings to DBpedia, which enabled us to obtain the matching for YAGO3.
Appendix B Evaluation methodology
B.1 Downstream tasks
Setting
For each dataset, we use scikit-learn’s Histogram-based Gradient Boosting Regression (resp. Classification) Tree [pedregosa2011scikit] as regression (resp. classification) estimator to predict the target value. The embeddings are the only features fed to the estimator, except for the US elections dataset for which we also include the political party. For embedding models outputting complex embeddings, such as RotatE, we simply concatenate real and imaginary parts before feeding them to the estimator.
The rows of the tables corresponding to entities not found in the knowledge graph are filled with NaNs as features for the estimator. This enables to compare the scores between different knowledge graphs (see Figure 2 and Figure 3) of different sizes to see the benefits obtained from embedding larger graphs, with better coverage of downstream entities (Figure 4).
Figure 5: Detailed results on real-world tables. The "Cumulative normalized mean cross-validation score" reported is obtained by summing the normalized mean cross-validation scores. For an evaluation dataset, 1 corresponds to the best R2 score across all models; as there are 4 evaluation datasets, the highest possible score for a model is 4 (getting a score of 4 means that the model beats every model on every evaluation dataset). SEPAL, PyTorch-BigGraph, DGL-KE, and NodePiece use DistMult as base model. Embedding computation times are provided on the right-hand side of the figure. Figure 7 extends this figure with other embedding models.
Figure 6: Detailed results on WikiDBs tables. Each black dot represents a downstream table. Red vertical lines indicate the mean score over all the tables. The methods appear in decreasing order of average score.
Metrics
The metric used for regression is the R2 score, defined as:
$R^{2}=1-\frac{\sum_{i=1}^{N}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}$
where $N$ is the number of samples (rows) in the target table, $y_{i}$ is the target value of sample $i$, $\hat{y}_{i}$ is the value predicted by the estimator, and $\bar{y}$ is the mean value of the target variable.
For classification, we use the weighted F1 score, defined as:
$F1_{\text{weighted}}=\sum_{i=1}^{K}\frac{n_{i}}{N}\cdot F1_{i}$
where $K$ is the number of classes, $n_{i}$ is the number of true instances for class $i$, $N=\sum_{i=1}^{K}n_{i}$ is the total number of samples, and $F1_{i}$ is the F1 score for class $i$.
To get the "Mean score (normalized)" reported on Figure 2 and Figure 3, we proceed as follows:
  1. 1.
    Mean cross-validation score: for each model [^2] We define “model” as the combination of a method (e.g. DistMult, DGL-KE, etc.) and a knowledge graph on which it is trained. and evaluation table, the scores (R2 or weighted F1, depending on the task) are averaged over 5 repeats of 5-fold cross-validations.
  2. 2.
    Normalized: for each evaluation table, we divide the scores of the different models by the score of the best-performing model on this table. This makes the scores more comparable between the different evaluation tables.
  3. 3.
    Average: for each model, we average its scores across every evaluation table. The highest possible score for a model is 1. Getting a score of 1 means that the model beats every other model on every evaluation table.
Validation/test split and hyperparameter tuning
We use 4 of the 42 WikiDBs tables as validation data—2 for regression and 2 for classification tasks (see Figure 4). The remaining 38 WikiDBs tables, along with the 4 real-world tables, are used exclusively for testing.
The validation tables are used to tune hyperparameters and select the best-performing configuration for each method and each knowledge graph. Unless otherwise specified, all reported results are obtained on the test tables using these optimized configurations.
B.2 Knowledge graph completion
We detail our experimental setup for the link prediction task:
Setting:
We evaluate models under the transductive setting: the missing links to be predicted connect entities already seen in the train graph. The task is to predict the tail entity of a triple, given its head and relation.
Stratification:
We randomly split each dataset into training (90%), validation (5%), and test (5%) subsets of triples. During stratification, we ensure that the train graph remains connected by moving as few triples as required from the validation/test sets to the training set.
Sampling:
Given the size of our datasets, sampling is required to keep link prediction tractable. For each evaluation triple, we sample 10,000 negative entities uniformly to produce 10,000 candidate negative triples by corrupting the positive.
Filtering:
For tractability reasons, we report unfiltered results: we do not remove triples already existing in the dataset (which may score higher than the test triple) from the candidates.
Ranking:
If several triples have the same score, we report realistic ranks (i.e., the expected ranking value over all permutations respecting the sort order; see PyKEEN documentation [ali2021pykeen]).
Metrics:
We use three standard metrics for link prediction: the mean reciprocal rank (MRR), hits at $k$ (for $k\in\{1,10,50\}$) and mean rank (MR). Given the rankings $r_{1},\ldots,r_{n}$ of the $n$ evaluation (validation or test) triples:
$\mathrm{MRR}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{r_{i}},$ $\mathrm{Hits@}k=\frac{1}{n}\sum_{i=1}^{n}1_{r_{i}\leq k},$ $\mathrm{MR}=\frac{1}{n}\sum_{i=1}^{n}r_{i}.$
B.3 Experimental setup
Baseline implementations
In our empirical study, we compare SEPAL to DistMult, NodePiece, PBG, DGL-KE, and FastRP. We use the PyKEEN [ali2021pykeen] implementation for DistMult and NodePiece, and the implementations provided by the authors for the others. PBG was trained on 20 CPU nodes, and DGL-KE was allocated 20 CPU nodes and 3 GPUs; both methods were run on a single machine. The version of NodePiece we use for datasets larger than Mini YAGO3 is the ablated version, where nodes are tokenized only from their relational context (otherwise, the method does not scale on our hardware). For PBG, DGL-KE, NodePiece, and FastRP we used the hyperparameters provided by the authors for datasets of similar sizes. SEPAL and DistMult’s hyperparameters were tuned on the validation sets presented in Section˜B.1 and Section˜B.2.
For all the baseline clustering algorithms, we used the implementations from the igraph package [csardi2013package] except for METIS, HDRF and Spectral Clustering. For METIS, we used the torch-sparse implementation, for Spectral Clustering, the scikit-learn [pedregosa2011scikit] implementation, and for HDRF, we used the C++ implementation from this repository.
Figure 7: Performance on real-world downstream tables.
Computer resources
For PBG and FastRP, experiments were carried out on a machine with 48 cores and 504 GB of RAM. DistMult, DGL-KE, NodePiece, and SEPAL were trained on Nvidia V100 GPUs with 32 GB of memory, and 20 CPU nodes with 252 GB of RAM. The clustering benchmark was run on a machine with 88 CPU nodes and 504 GB of RAM.
Table 6: Results for link prediction. Best in bold, second underlined.
YAGO3 YAGO4.5 YAGO4.5+T YAGO4 YAGO4+T Freebase WikiKG90Mv2 Average
a. MRR
DistMult 0.8049 - - - - - - -
NodePiece 0.2596 0.4456 - - - - - -
PBG 0.5581 0.5539 0.5688 0.6406 0.6224 0.7389 0.6325 0.6165
DGL-KE 0.7284 0.6200 0.6469 0.2372 0.2570 0.3017 0.3202 0.4445
SEPAL 0.6501 0.5537 0.5646 0.4726 0.477 0.5378 0.5291 0.5407
b. Hits@1
DistMult 0.7400 - - - - - -
NodePiece 0.1735 0.3449 - - - - -
PBG 0.5000 0.4939 0.4977 0.5494 0.5416 0.7015 0.5568 0.5487
DGL-KE 0.6663 0.5511 0.5733 0.1642 0.1892 0.2498 0.2416 0.3765
SEPAL 0.5412 0.4913 0.4922 0.3746 0.3755 0.4824 0.4502 0.4582
c. Hits@10
DistMult 0.9059 - - - - - - -
NodePiece 0.4388 0.6379 - - - - - -
PBG 0.6562 0.6642 0.7021 0.803 0.7662 0.8053 0.7693 0.7380
DGL-KE 0.8293 0.7446 0.7821 0.3786 0.3842 0.3964 0.4694 0.5692
SEPAL 0.8394 0.6650 0.6871 0.6573 0.6778 0.6398 0.6739 0.6915
d. Hits@50
DistMult 0.9504 - - - - - - -
NodePiece 0.7358 0.8112 - - - - - -
PBG 0.7259 0.7734 0.8136 0.8891 0.8514 0.8541 0.8531 0.8229
DGL-KE 0.8873 0.8171 0.8748 0.6037 0.5968 0.5547 0.654 0.7126
SEPAL 0.9204 0.7698 0.7786 0.7661 0.8068 0.7476 0.7805 0.7957
e. MR
DistMult 64.19 - - - - - - -
NodePiece 154.7 263.2 - - - - - -
PBG 820.9 408.0 300.3 117.1 203.5 243.4 227.3 331.5
DGL-KE 187.7 624.1 219.9 224.9 271.5 227.0 186 277.3
SEPAL 95.87 270.4 206.0 553.0 363.2 357.3 365.5 315.9
Appendix C Additional evaluation of SEPAL
C.1 Table-level downstream results on real-world tables
Figure 5 shows the detailed prediction performance on the downstream tasks. SEPAL not only scales well to very large graphs (computing times markedly smaller than Pytorch-BigGraph), but also creates more valuable node features for downstream tasks.
C.2 Table-level downstream results on WikiDBs tables
Figure 6 shows the detailed downstream tasks results for the WikiDBs test tables. For regression, SEPAL is the best performer on average for all of the 7 knowledge graphs. For classification, SEPAL is the best method on 6 of 7 knowledge graphs (narrowly beaten by PBG on YAGO4.5). Interestingly, FastRP, despite being very simple and not even accounting for relations, is a strong performer and beats more sophisticated methods like PBG and DGL-KE on many tasks. This is consistent with our queriability analysis in section 4 concluding that global methods, such as FastRP and SEPAL, are more suitable for downstream tasks, than local methods (such as PBG and DGL-KE).
For some knowledge graphs (especially Freebase and YAGO3), we can see two modes in the scores distribution, corresponding to the tables with low and high coverages (Figure 4).
C.3 SEPAL combined with more embedding models
Figure 7 extends the results of Figure 5 by adding TransE and RotatE, alone and combined with SEPAL, as well as ablation studies results of SEPAL combined with METIS or ablated from BLOCS. These results show that SEPAL systematically improves upon its base model, whether it be DistMult RotatE or TransE. For a fair comparison, we ran RotatE with embedding dimension $d=50$, as it outputs complex embeddings having twice as many parameters. For other models, we use $d=100$.
C.4 More baselines on the Freebase dataset
The Freebase dataset has been used in several previous works. To further validate SEPAL’s performance, we compare it to additional baselines from the large-scale KGE literature: GraSH [kochsiek2022start], an efficient hyperparameter optimization framework for large-scale KGEs, and SMORE [ren2022smore], a scalable KGE method supporting single-GPU training and multi-hop reasoning.
We train both of these baselines with DistMult as the base embedding method, on Freebase, using the authors’ released configurations for this dataset. Table 7 reports the normalized scores on four real-world regression tasks, along with total training time.
Table 7: Comparison to additional baselines on Freebase. Normalized mean cross-validation scores on real-world downstream tasks, along with total training time. SEPAL outperforms all baselines while being the fastest to train.
Method Housing prices Movie revenues US accidents US elections Time
PBG 0.513 0.723 0.742 0.957 33h 42m
DGL-KE 0.445 0.610 0.686 0.962 10h 50m
SMORE 0.160 0.332 0.410 0.926 6h 15m
GraSH 0.601 0.862 0.810 0.961 32h 33m
SEPAL 0.868 0.880 0.953 1.000 5h 58m
We trained SMORE for 1 million iterations on one GPU, following the authors’ configuration. Its relatively low performance here suggests that longer training could improve results, but under a 6-hour budget, SEPAL is substantially better.
GraSH optimizes hyperparameters for link prediction using successive halvings to discard unpromising configurations at low cost. Results show that, after 33 hours of hyperparameter search on one GPU, GraSH produces better embeddings than other baselines. However, SEPAL remains the best performer on all tasks, and also the fastest method.
C.5 Evaluation on prior benchmark
ruffinelli2024beyond propose a benchmark for evaluating KGE methods on downstream classification and regression. It includes several KGE baselines, with varying training procedures, and the graph neural network KE-GCN for entity classification. For each knowledge graph (FB15k-237, YAGO3-10, Wikidata5M), they evaluate embeddings on downstream tasks created from entity attributes of the knowledge graph.
We put the evaluation on this benchmark in appendix because our goal in this paper is to evaluate external, real-world tasks (Figure 2) that are independent of a specific knowledge graph, to compare the benefits of diverse knowledge graphs. In contrast, the Ruffinelli et al. tasks are created artificially and associated with specific knowledge graphs, that are orders of magnitude smaller than those of our main study.
However, evaluating SEPAL on these datasets complements our main experiments and enables direct comparison with prior work. Thus, we used 128-dimensional embeddings to match one of the dimensions in Ruffinelli et al.’s hyperparameter search space. We also used the authors’ released evaluation script for comparable results. For each knowledge graph, Table 8 reports:
  1. 1.
    The best KGE method for classification from ruffinelli2024beyond;
  2. 2.
    The best KGE method for regression from ruffinelli2024beyond;
  3. 3.
    KE-GCN results from ruffinelli2024beyond;
  4. 4.
    SEPAL results.
Table 8: Evaluation on Ruffinelli et al.’s benchmark. Classification and regression results on FB15k-237 (14k entities, 272k triples), YAGO3-10 (123k entities, 1M triples), and Wikidata5M (4.8M entities, 21M triples). Best results for each task are in bold.
Dataset Method Classification weighted F1 $\uparrow$ Regression RSE $\downarrow$
FB15k-237 ComplEx (MTT) 0.858 0.394
RotatE (MTT) 0.890 0.573
KE-GCN 0.829 0.501
SEPAL 0.853 0.492
YAGO3-10 DistMult (MTT) 0.746 0.472
TransE (MTT) 0.723 0.441
KE-GCN 0.700 0.398
SEPAL 0.762 0.386
Wikidata5M TransE (STD) 0.596
SEPAL 0.568
The results show that, on YAGO3-10 and Wikidata5M, SEPAL achieves the best performance for both regression and classification, consistent with our main results on much larger knowledge graphs. On FB15k-237, SEPAL has weaker results, but this may be due to the dataset size (FB15k-237 has 14k entities, 185 to 6,500 times smaller than the graphs considered in our main evaluation). For these tiny graphs, which fall outside SEPAL’s intended scope, SEPAL may not be adapted because the core becomes too small to learn good representations for the relations and core entities.
C.6 Evaluation on knowledge graph completion
C.6.1 Link prediction results
Table 6 provides the link prediction results for the different metrics. It shows that, depending on the dataset, SEPAL is competitive with existing methods (DGL-KE, PBG) or not. However, none of the methods is consistently better than the others for all datasets. Following the analysis in section 4, these results are expected given that SEPAL does not enforce local contrast between positive and negative triples through contrastive learning with negative sampling, like other KGE methods do, but rather focuses on global consistency of the embeddings. Link prediction is, by essence, a local task, asking to discriminate efficiently between positives and negatives. For this purpose, it seems that the negative sampling, absent from SEPAL’s propagation, plays a crucial role.
Figure 8: Critical difference diagrams on link prediction metrics. Statistical tests show no significant difference between methods at significance level $\alpha=0.05$.
Figure 9: Downstream task performance against link prediction performance, on validation sets. Linear regression, with a 95% confidence interval. We plot the performance of models that share the same hyperparameters on the train graph (used for link prediction) and the full graph (used for downstream tasks).
Nevertheless, Figure 8 shows that a Friedman test followed by Conover’s post-hoc analysis [conover1979multiple, conover1999practical] reveal no statistically significant differences (at significance level $\alpha=0.05$) among the three methods that scale to all the knowledge graphs (SEPAL, PBG, and DGL-KE), as indicated by the critical difference diagrams where all methods are connected by a black line.
From a hyperparameter perspective, contrary to downstream tasks, the hybrid core selection strategy (with both node and relation sampling) yields better results than its simpler degree-based counterpart. This highlights different trade-offs between downstream tasks and link prediction. For link prediction, good relational coverage seems to count most, whereas for downstream tasks, the core density matters most (Figure 20).
C.6.2 Downstream and link prediction performance weakly correlate
Figure 9 shows that embeddings performing well on downstream tasks do not necessarily perform well on link prediction, and vice versa. There is only a small positive correlation between these two performances: $R^{2}\sim 0.1$.
Appendix D Theoretical analysis
Notations
We use the following notations:
  • $\boldsymbol{\Theta}^{(t)}\in\mathbb{R}^{n\times d}$ is the embedding matrix at step $t$, where each row is the embedding of an entity. Without loss of generality, we assume that the entities are ordered core first, then outer, so we can write $\boldsymbol{\Theta}^{(t)}=\begin{bmatrix}\boldsymbol{\Theta}_{c}\\ \boldsymbol{\Theta}_{o}^{(t)}\end{bmatrix}$ with $\boldsymbol{\Theta}_{c}\in\mathbb{R}^{n_{c}\times d}$ the embedding matrix of core (fixed) entities, and $\boldsymbol{\Theta}_{o}^{(t)}\in\mathbb{R}^{n_{o}\times d}$ the embedding matrix of outer (updated) entities at step $t$. $n_{c}$ and $n_{o}$ denote the number of core and outer entities, respectively, and $n_{c}+n_{o}=n$ is the total number of entities.
  • $\boldsymbol{w}_{r}\in\mathbb{R}^{d}$: embedding of relation $r\in\mathcal{R}$ (fixed).
  • $\boldsymbol{x}^{(t)}=\operatorname{vec}(\boldsymbol{\Theta}^{(t)})=\left[\boldsymbol{\Theta}^{(t)}_{1,1},\ldots,\boldsymbol{\Theta}^{(t)}_{n,1},\boldsymbol{\Theta}^{(t)}_{1,2},\ldots,\boldsymbol{\Theta}^{(t)}_{n,2},\ldots,\boldsymbol{\Theta}^{(t)}_{1,d},\ldots,\boldsymbol{\Theta}^{(t)}_{n,d}\right]^{\top}\in\mathbb{R}^{nd}$: vectorization of the embedding matrix.
  • $\boldsymbol{P}\in\mathbb{R}^{nd\times nd}$: global linear propagation matrix.
  • $\boldsymbol{Q}\in\mathbb{R}^{nd\times nd}$: permutation matrix to reorder $\boldsymbol{x}$ into core and outer blocks.
  • $\boldsymbol{y}^{(t)}=\boldsymbol{Q}\boldsymbol{x}^{(t)}=\left[\boldsymbol{\theta}_{1}^{\top};\ldots;\boldsymbol{\theta}_{n}^{\top}\right]^{\top}\in\mathbb{R}^{nd}$: reordered vector of embeddings. $\boldsymbol{y}^{(t)}$ can also be written as $\boldsymbol{y}^{(t)}=\begin{bmatrix}\boldsymbol{y}_{c}\\ \boldsymbol{y}_{o}^{(t)}\end{bmatrix}$.
  • $\boldsymbol{M}=\boldsymbol{Q}(\boldsymbol{I}+\alpha\boldsymbol{P})\boldsymbol{Q}^{-1}\in\mathbb{R}^{nd\times nd}$: reordered update matrix. $\boldsymbol{M}$ can be written by block $\boldsymbol{M}=\begin{bmatrix}\boldsymbol{M}_{cc}&\boldsymbol{M}_{co}\\ \boldsymbol{M}_{oc}&\boldsymbol{M}_{oo}\end{bmatrix}$ where $\boldsymbol{M}_{oo}$ and $\boldsymbol{M}_{oc}$ are submatrices representing outer-to-outer and core-to-outer influences.
D.1 Analysis of SEPAL’s dynamic and analogies to eigenvalue problems
This section presents a theoretical analysis of the SEPAL propagation algorithm. We provide a series of intuitive and structural analogies to classic iterative methods in numerical linear algebra. Our goal is to contextualize the algorithm’s behavior under various assumptions and shed light on its dynamic properties. The analysis is carried out in the case of DistMult, which simplifies the propagation rule due to its element-wise multiplication structure.
SEPAL as power iteration (no normalization, no boundary conditions)
We begin by analyzing the case without normalization or fixed embeddings. In this setting, with the vectorized embedding matrix $\boldsymbol{x}^{(t)}\in\mathbb{R}^{nd}$, the propagation equation (Equation 6) becomes:
$\boldsymbol{x}^{(t+1)}=(\boldsymbol{I}+\alpha\boldsymbol{P})\boldsymbol{x}^{(t)},$ (10)
where $\boldsymbol{P}\in\mathbb{R}^{nd\times nd}$ encodes the linear update based on the knowledge graph structure and DistMult composition rule. The matrix $\boldsymbol{P}$ is block diagonal:
$\boldsymbol{P}=\begin{bmatrix}\boldsymbol{P}^{(1)}&\boldsymbol{0}&\cdots&\boldsymbol{0}\\ \boldsymbol{0}&\boldsymbol{P}^{(2)}&\cdots&\boldsymbol{0}\\ \vdots&\vdots&\ddots&\vdots\\ \boldsymbol{0}&\boldsymbol{0}&\cdots&\boldsymbol{P}^{(d)}\end{bmatrix}\in\mathbb{R}^{nd\times nd},$ (11)
where each block $P^{(k)}$ corresponds to the $k$-th embedding dimension and has:
$\boldsymbol{P}^{(k)}_{u,v}=\sum_{(v,r,u)\in\mathcal{K}}[\boldsymbol{w}_{r}]_{k}.$ (12)
$\boldsymbol{P}^{(k)}$ can be seen as a weighted adjacency matrix of the graph, whose weights are the $k$-th coefficients of the relation embeddings of the corresponding edges. Here, each $(v,r,u)\in\mathcal{K}$ contributes a rank-1 update to $\boldsymbol{P}$ based on $\boldsymbol{w}_{r}$.
Therefore, in this setting, the problem is separable with respect to the dimensions, so each dimension can be studied independently.
The recurrence in Equation 10 defines a classical power iteration. In general, it diverges unless the spectral radius $\rho(\boldsymbol{\boldsymbol{I}+\alpha\boldsymbol{P}})<1$. In our setting, norms can grow arbitrarily because $\boldsymbol{P}$ contains non-normalized adjacency submatrices whose eigenvalues are only bounded by the maximum degree of the graph. In practice, the normalization (studied below) prevents the algorithm from diverging.
With boundary conditions: non-homogeneous recurrence
Now, we consider the SEPAL’s setting where core entity embeddings are fixed and only outer embeddings evolve, still without normalization. We use the reordered vector of embeddings $\boldsymbol{y}^{(t)}=\boldsymbol{Q}\boldsymbol{x}^{(t)}\in\mathbb{R}^{nd}$, which can be written as follows:
$\boldsymbol{y}^{(t)}=\begin{bmatrix}\boldsymbol{\theta}_{1}^{(t)}\\ \boldsymbol{\theta}_{2}^{(t)}\\ \vdots\\ \boldsymbol{\theta}_{n}^{(t)}\end{bmatrix}=\begin{bmatrix}\boldsymbol{y}_{c}\\ \boldsymbol{y}_{o}^{(t)}\end{bmatrix}$, (13)
where $\boldsymbol{\theta}_{u}^{(t)}$ is the embedding of entity $u$. The reordered update matrix $\boldsymbol{M}=Q(\boldsymbol{I}+\alpha\boldsymbol{P})Q^{-1}$ has the following block structure:
$\boldsymbol{M}=\begin{bmatrix}\boldsymbol{M}_{11}&\cdots&\boldsymbol{M}_{1n}\\ \vdots&\ddots&\vdots\\ \boldsymbol{M}_{n1}&\cdots&\boldsymbol{M}_{nn}\end{bmatrix}\in\mathbb{R}^{nd\times nd},$ (14)
where each block $\boldsymbol{M}_{uv}\in\mathbb{R}^{d\times d}$ encodes how the embedding of entity $v$ contributes to the update of entity $u$. In the case of the DistMult scoring function, this block is diagonal and takes the form:
$\boldsymbol{M}_{uv}=\begin{cases}\sum_{(v,r,u)\in\boldsymbol{\mathcal{K}}}\alpha\cdot\operatorname{diag}(\boldsymbol{w}_{r})&\text{if $u\neq v$,}\\ \boldsymbol{I}_{d}+\sum_{(v,r,u)\in\boldsymbol{\mathcal{K}}}\alpha\cdot\operatorname{diag}(\boldsymbol{w}_{r})&\text{otherwise,}\end{cases}$ (15)
where $\operatorname{diag}(\boldsymbol{w}_{r})$ is the diagonal matrix with the relation embedding $\boldsymbol{w}_{r}\in\mathbb{R}^{d}$ on the diagonal.
Thus, $\boldsymbol{M}$ is a sparse block matrix, with each non-zero block being diagonal, and it linearly propagates information across entity embeddings via dimension-wise interactions determined by the DistMult model. Grouping core and outer entities together, we can also write $\boldsymbol{M}=\begin{bmatrix}\boldsymbol{M}_{cc}&\boldsymbol{M}_{co}\\ \boldsymbol{M}_{oc}&\boldsymbol{M}_{oo}\end{bmatrix}$.
This allows us to rewrite the propagation equation to account for the boundary conditions. We obtain:
$\boldsymbol{y}_{o}^{(t+1)}=\boldsymbol{M}_{oo}\boldsymbol{y}_{o}^{(t)}+\boldsymbol{M}_{oc}\boldsymbol{y}_{c},$ (16)
where $\boldsymbol{M}_{oo}$ represents signal propagation between outer nodes, and $\boldsymbol{M}_{oc}$ encodes injection from the core nodes.
This is a non-homogeneous linear recurrence. If $\rho(\boldsymbol{M}_{oo})\geq 1$, the outer embeddings diverge in norm. Nonetheless, the structure is analogous to forced linear systems such as:
$\boldsymbol{y}_{t+1}=\boldsymbol{A}\boldsymbol{y}_{t}+\boldsymbol{b},$
where the long-term behavior is driven by the balance between eigenvalues of $\boldsymbol{A}$ and direction of $\boldsymbol{b}$.
Normalization and Arnoldi-type analogy
SEPAL applies $\ell_{2}$ normalization after each update:
$\boldsymbol{\theta}_{u}^{(t+1)}=\frac{\boldsymbol{\theta}_{u}^{(t)}+\alpha\boldsymbol{a}_{u}^{(t+1)}}{\|\boldsymbol{\theta}_{u}^{(t)}+\alpha\boldsymbol{a}_{u}^{(t+1)}\|_{2}},$ (17)
which constrains every embedding to the unit sphere. This couples dimensions and prevents simple linear analysis.
However, the recurrence
$\boldsymbol{y}_{o}^{(t+1)}=\boldsymbol{M}_{oo}\boldsymbol{y}_{o}^{(t)}+\boldsymbol{M}_{oc}\boldsymbol{y}_{c},$ (18)
followed by blockwise normalization of $\boldsymbol{y}^{(t)}$ (with a $\ell_{2,\infty}$ mixed norm), shares structural similarities with the Arnoldi iteration [arnoldi1951principle]:
  • Successive embeddings span a Krylov subspace: after iteration $t$, without normalization, $\boldsymbol{y}_{o}^{(t)}$ belongs to $\mathcal{K}_{t}(\boldsymbol{A},\boldsymbol{b})=\operatorname{span}\,\{\boldsymbol{b},\boldsymbol{A}\boldsymbol{b},\boldsymbol{A}^{2}\boldsymbol{b},\ldots,\boldsymbol{A}^{t-1}\boldsymbol{b}\}$, with $\boldsymbol{A}=\boldsymbol{M}_{oo}$ and $\boldsymbol{b}=\boldsymbol{M}_{oc}\boldsymbol{y}_{c}$, given that $\boldsymbol{y}_{o}^{(0)}=\boldsymbol{0}$.
  • Core embeddings define the forcing direction ($\boldsymbol{b}=\boldsymbol{M}_{oc}\boldsymbol{y}_{c}$).
  • Normalization serves as a regularizer, preventing divergence in norm.
The Arnoldi iteration is used to compute numerical approximations of the eigenvectors of general matrices, for instance, in ARPACK [lehoucq1998arpack]. This analogy suggests that the direction of outer embeddings stabilizes over time and aligns with a form of dominant generalized eigenvector of the propagation operator $\boldsymbol{M}$, conditioned on the core. Therefore, the embeddings produced by SEPAL’s propagation encapsulate global structural information on the knowledge graph.
D.2 Formal proof of Proposition˜4.1
Gradient descent on $\mathcal{E}$ updates the embedding parameters $\boldsymbol{\theta}_{u}$ of an entity $u$ with
$\boldsymbol{\theta}_{u}^{(t+1)}=\boldsymbol{\theta}_{u}^{(t)}-\eta\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}_{u}^{(t)}}$ (19)
where $\eta$ is the learning rate.
The mini-batch gradient satisfies
$\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}_{u}^{(t)}}$ $=-\sum_{(h,r,t)\in\mathcal{B}}\frac{\partial\left\langle\boldsymbol{\theta}_{t}^{(t)},\phi(\boldsymbol{\theta}_{h}^{(t)},\boldsymbol{w}_{r})\right\rangle}{\partial\boldsymbol{\theta}_{u}^{(t)}}$
$=-\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ h=u\end{subarray}}\frac{\partial\left\langle\boldsymbol{\theta}_{t}^{(t)},\phi(\boldsymbol{\theta}_{u}^{(t)},\boldsymbol{w}_{r})\right\rangle}{\partial\boldsymbol{\theta}_{u}^{(t)}}-\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ t=u\end{subarray}}\frac{\partial\left\langle\boldsymbol{\theta}_{u}^{(t)},\phi(\boldsymbol{\theta}_{h}^{(t)},\boldsymbol{w}_{r})\right\rangle}{\partial\boldsymbol{\theta}_{u}^{(t)}},$
where $\mathcal{B}$ is the mini-batch. Note that the previous identity is still true if the knowledge graph contains self-loops, due to DistMult’s scoring function being a product. Plugging the DistMult relational operator function $\phi(\boldsymbol{\theta}_{h},\boldsymbol{w}_{r})=\boldsymbol{\theta}_{h}\odot\boldsymbol{w}_{r}$ in the previous equation gives
$\frac{\partial\mathcal{E}}{\partial\boldsymbol{\theta}_{u}^{(t)}}$ $=-\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ h=u\end{subarray}}\frac{\partial{\boldsymbol{\theta}_{u}^{(t)}}^{\top}(\boldsymbol{w}_{r}\odot\boldsymbol{\theta}_{t}^{(t)})}{\partial\boldsymbol{\theta}_{u}^{(t)}}-\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ t=u\end{subarray}}\frac{\partial{\boldsymbol{\theta}_{h}^{(t)}}^{\top}(\boldsymbol{w}_{r}\odot\boldsymbol{\theta}_{u}^{(t)})}{\partial\boldsymbol{\theta}_{u}^{(t)}}$
$=-\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ h=u\end{subarray}}\boldsymbol{w}_{r}\odot\boldsymbol{\theta}_{t}^{(t)}-\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ t=u\end{subarray}}\boldsymbol{\theta}_{h}^{(t)}\odot\boldsymbol{w}_{r}$
$=-\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ h=u\end{subarray}}\boldsymbol{w}_{r}\odot\boldsymbol{\theta}_{t}^{(t)}-\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ t=u\end{subarray}}\phi(\boldsymbol{\theta}_{h}^{(t)},\boldsymbol{w}_{r}).$
Therefore, going back to Equation 19, we get that
$\underbrace{\boldsymbol{\theta}_{u}^{(t+1)}=\boldsymbol{\theta}_{u}^{(t)}+\eta\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ t=u\end{subarray}}\phi(\boldsymbol{\theta}_{h}^{(t)},\boldsymbol{w}_{r})}_{\text{embedding propagation update for }\eta=\alpha\text{ and }\mathcal{B}=\mathcal{S}\cup\mathcal{C}}+\eta\sum_{\begin{subarray}{c}(h,r,t)\in\mathcal{B}\\ h=u\end{subarray}}\boldsymbol{w}_{r}\odot\boldsymbol{\theta}_{t}^{(t)}.$ (20)
We can see that the embedding propagation update only differs by a term that corresponds to the message passing from the tails to the heads. We did not include this term in our message-passing framework because we wanted SEPAL to adapt to any model whose scoring function has the form given by Equation 2. In practice, the propagation direction tail to head is already handled by the addition of inverse relations.
Therefore, a parallel can be drawn between: 1) the outer subgraphs and mini-batches, 2) the number of propagation steps $T$ and the number of epochs, 3) the hyperparameter $\alpha$ and the learning rate.
After each gradient step, we normalize the entity embeddings to enforce the unit norm constraint. This procedure corresponds to projected gradient descent on the sphere [bertsekas1997nonlinear], where each update is followed by a projection (via $\ell^{2}$ normalization) back onto the feasible set. The energy function $\mathcal{E}$ (Equation 7) is composed of inner products and element-wise multiplications of smooth functions, thus it is smooth, and its gradient is Lipschitz continuous on the unit sphere. Under these conditions, the algorithm is guaranteed to converge to a stationary point of the constrained optimization problem [bertsekas1997nonlinear, Proposition 2.3.2]. The limit points of the optimization thus satisfy the first-order optimality conditions on the sphere.
Algorithm 1 BLOCS
  Input: Graph $\mathcal{G}=(V,E)$ with nodes $V$ and edges $E$, hyperparameters $h$ and $m$
  Output: List of overlapping connected subgraphs
  $\mathbb{S}\leftarrow\emptyset$ $\triangleright$ list of subgraphs
  $U\leftarrow V$ $\triangleright$ set of unassigned nodes
  Step 1: Create subgraphs from super-spreaders’ neighbors
  for each node $v\in V$ do
   if deg(v)>0.2mdeg(v)>0.2\,m then
    $\mathbb{S},U\leftarrow\texttt{SplitNeighbors}(v,\,\text{max\_size}=0.2\,m)$
   end if
  end for
  Step 2: Assign nodes to subgraphs by diffusion
  while |U|>(1h)|V||U|>(1-h)|V| do
   $k\leftarrow 0$    ;    $\mathcal{S}_{0}\leftarrow\{\operatorname*{arg\,max}_{v\in U}deg(v)\}$ $\triangleright$ start with unassigned node $v$ with highest degree
   while $|\mathcal{S}_{k}|<0.8\,m$ do
    $\mathcal{S}_{k+1}\leftarrow\texttt{Diffuse}(\mathcal{S}_{k})$    ;    $k\leftarrow k+1$
   end while
   Append $\mathcal{S}_{k-1}$ to $\mathbb{S}$, and update $U$ $\triangleright$ $\mathcal{S}_{k-1}$ is the last subgraph smaller than 0.8m0.8\,m
  end while
  Step 3: Merge small overlapping subgraphs
  $\mathbb{S},U\leftarrow\texttt{MergeSmallSubgraphs}(\mathbb{S},\,\text{min\_size}=m/2)$
  Step 4: Dilation and diffusion until all entities are assigned
  $p\leftarrow 0\triangleright$ create new subgraphs by diffusion every 5 rounds, to tackle long chains
  while $|U|>0$ do
   if 5 divides $p$ and $p>0$ then
    $i\leftarrow 0$
    repeat
     $k\leftarrow 0$    ;    $\mathcal{S}_{0}\leftarrow\{\operatorname*{arg\,max}_{v\in U}deg(v)\}$
     while $|\mathcal{S}_{k}|<0.8\,m$ do
      $\mathcal{S}_{k+1}\leftarrow\texttt{Diffuse}(\mathcal{S}_{k})$    ;    $k\leftarrow k+1$
     end while
     Append $\mathcal{S}_{k-1}$ to $\mathbb{S}$, and update $U$    ;    $i\leftarrow i+1$
    until $i=10$
   end if
   $\mathbb{S}\leftarrow\texttt{Dilate}(\mathbb{S})$    ;    $p\leftarrow p+1$
  end while
  Step 5: Merge small overlapping subgraphs again
  $\mathbb{S},U\leftarrow\texttt{SystematicMerge}(\mathbb{S},\,\text{min\_size}=0.4\,m)$
  Step 6: Split subgraphs larger than $m$
  $\mathbb{S},U\leftarrow\texttt{SplitLargeSubgraphs}(\mathbb{S},\,\text{max\_size}=m)$
  $\mathbb{S},U\leftarrow\texttt{MergeSmallSubgraphs}(\mathbb{S},\,\text{min\_size}=m/2)$
  Return: $\mathbb{S}$, the set of overlapping subgraphs covering $\mathcal{G}$
Appendix E Presentation and analysis of BLOCS
E.1 Prior work on graph partitioning
Scaling up computation on graph, for graph embedding or more generally, often relies on breaking down graphs in subgraphs. METIS [karypis1997metis], a greedy node-merging algorithm, is a popular solution. A variety of algorithms have also been developed to detect “communities”, groups of nodes more connected together, often with applications on social networks: Spectral Clustering (SC) [shi2000normalized], the Leading Eigenvector (LE) method [newman2006finding], the Label Propagation Algorithm (LPA) [raghavan2007near], the Louvain method [blondel2008fast], the Infomap method [rosvall2008maps], and the Leiden method [traag2019louvain] which guarantees connected communities. LDG [stanton2012streaming] and FENNEL [tsourakakis2014fennel] are streaming algorithms for very large graphs. Some algorithms have also been specifically tailored for power-law graphs: DBH [xie2014distributed] leverages the skewed degree distributions to reduce the communication costs, HDRF [petroni2015hdrf] is a streaming partitioning method that replicates high-degree nodes first, and Ginger [chen2019powerlyra] is a hybrid-cut algorithm that uses heuristics for more efficient partitioning on skewed graphs.
E.2 Detailed algorithm and pseudocode
Figure 10: Scalability of partitioning methods. Memory usage and time for 8 knowledge graphs of various sizes. The “CPU RAM capacity” dashed line represents the CPU RAM of the machine we used to run SEPAL (we used a different machine with more RAM to run this benchmark, see Section˜B.3). Partitioning methods going beyond this limit thus cannot be combined with SEPAL on our hardware. BLOCS is the only method to scale up to WikiKG90Mv2, a knowledge graph with 601M triples. Leiden and METIS both caused memory errors on WikiKG90Mv2, while HDRF was too long for graphs larger than YAGO4 (our time limit was set to 333 minutes).
Figure 11: Quality of partitioning methods. Maximum number of connected components in one partition, ratio between the largest and smallest partition sizes, and size (number of entities) of the largest partition produced for knowledge graphs of various sizes. The “GPU capacity” dashed line represents the typical number of entities that can be loaded onto the GPU before causing a memory error. Methods producing partitions larger than this cannot be combined with SEPAL, since the partition’s embeddings must fit in GPU memory. BLOCS and Leiden are the only methods to consistently return connected partitions (requirement R1). BLOCS, METIS, and HDRF are the only methods to control the partition size (requirement R2).
Algorithm˜1 describes BLOCS pseudocode. Below, we provide more details on subparts of the algorithm.
The function SplitNeighbors is designed to manage high-degree nodes by distributing their neighboring entities into smaller subgraphs. For each node whose degree exceeds a certain threshold, it groups its neighbors into multiple subgraphs smaller than this threshold. These new subgraphs also include the original high-degree node to maintain connectedness. By assigning the neighbors of the very high-degree nodes first, BLOCS ensures that subgraphs grow more progressively in the subsequent diffusion step.
The function MergeSmallSubgraphs takes a list of subgraphs and a minimum size threshold as input. It identifies subgraphs that are smaller than this given minimum size and merges them into larger subgraphs if they share nodes and if the size of their union remains below the maximum size $m$.
The function SystematicMerge is very similar to MergeSmallSubgraphs, but it does not check that the resulting subgraphs are smaller than $m$. Its objective is to eliminate all the subgraphs whose size is smaller than a given threshold (set to 0.4m0.4\,m in Algorithm˜1). The subgraphs produced that are larger than $m$ are then handled by the function SplitLargeSubgraphs.
The function SplitLargeSubgraphs processes the list of subgraphs to break down overly large subgraphs while preserving connectivity. The function iterates over the subgraphs having more than $m$ nodes, subtracts the core, and computes the connected components. Then, it creates new subgraphs by grouping these connected components until the size limit $m$ is reached. As a result, outer subgraphs can be disconnected at this stage, but merging them with the core ensures their connectedness during embedding propagation.
E.3 Benchmarking BLOCS against partitioning methods
First, we compare BLOCS to other graph partitioning, clustering, and community detection methods. Figure 10 and Figure 11 report empirical evaluation on eight knowledge graphs. BLOCS, METIS, and Leiden are the only approaches that scale to the largest knowledge graphs. Others fail due to excessive runtimes –our limit was set to $2\cdot 10^{4}$ seconds. Compared to METIS, BLOCS is more efficient in terms of RAM usage while having similar computation times (Figure 10). Experimental results also show that classic partitioning methods fail to meet the connectedness and size requirements. Indeed, knowledge graphs are prone to yield disconnected partitions due to their scale-free nature: they contain very high-degree nodes. Such a node is hard to allocate to a single subgraph, and subgraphs without it often explode into multiple connected components. Our choice of overlapping subgraphs avoids this problem.
Classic methods do not meet the requirements of SEPAL
Here, we provide qualitative observations on the partitions produced by the different methods. We explain why they fail to meet our specific requirements.
METIS
is based on a multilevel recursive bisection approach, which coarsens the graph, partitions it, and then refines the partitions. It produces partitions with the same sizes; however, due to the graph structure, they often explode into multiple connected components, which is detrimental to the embedding propagation (see Section˜E.3).
Louvain
is based on modularity optimization. It outputs highly imbalanced communities, often with one community containing almost all the nodes and a few very small communities. This imbalance is incompatible with our approach, which requires strict control over the size of the subgraphs so that their embedding fits in GPU memory. Moreover, some of the communities are disconnected. On the two largest graphs, YAGO4 + taxonomy and Freebase, Louvain exceeds the time limit ($2\cdot 10^{4}$ seconds).
Leiden
modifies Louvain to guarantee connected communities and more stable outputs. It has very good scaling capabilities (Figure 10) but shares with Louvain the issue of producing highly-imbalanced communities.
LE
is a recursive algorithm that splits nodes based on the sign of their corresponding coefficient in the leading eigenvector of the modularity matrix. If these signs are all the same, the algorithm does not split the network further. Experimentally, LE returns only one partition (containing all the nodes) for YAGO3, YAGO4, YAGO4 + taxonomy, and Freebase. For YAGO4.5 + taxonomy, it hits our pre-set time limit ($2\cdot 10^{4}$ seconds). Therefore, we only report its performance for Mini YAGO3 and YAGO4.5, for which it outputs 2 and 4 partitions, respectively. It is important to note that the more communities it returns, the longer it takes to run because the algorithm proceeds recursively.
Infomap
uses random walks and information theory to group nodes into communities. Experimentally, it produces a lot of small communities with no connectedness guarantee. Additionally, it is too slow to be used on large graphs.
LPA
propagates labels across the network iteratively, allowing densely connected nodes to form communities. Similarly to Louvain and Leiden, the downside is that it does not control the size of the detected communities. It is also too slow to run on the largest graphs.
SC
uses the smallest eigenvectors of the graph Laplacian to transform the graph into a low-dimensional space and then applies k-means to group nodes together. However, the expensive eigenvector computation is a bottleneck that does not allow this approach to be used on huge graphs.
HDRF
is a streaming algorithm that produces balanced edge partitions (vertex cut) and minimizes the replication factor. However, it produces disconnected partitions, and it is too slow to run on the largest graphs.
Figure 12: Ablation study: replacing BLOCS with METIS. Normalized R2 scores (same as Figure 5) aggregated across evaluation datasets (movie revenues, US accidents, US elections, housing prices) for SEPAL with BLOCS and METIS are plotted against CPU RAM usage. BLOCS necessitates significantly less memory than METIS. We were not able to run SEPAL + METIS on knowledge graphs larger than YAGO4.5, hitting CPU RAM limits during the partitioning stage.
Therefore, none of the above methods readily produces subgraphs suitable for SEPAL. Indeed Figure 11 shows that BLOCS is the only method that returns subgraphs that are both connected and bounded in size, while being competitive in terms of scalability (Figure 10).
BLOCS cannot be replaced with METIS
To demonstrate the benefits of BLOCS over existing methods, we try to replace BLOCS with METIS in our framework. The results are presented in Figure 12.
Two important points differentiate these methods:
  1. 1.
    Contrary to BLOCS, METIS outputs disconnected partitions (see Figure 11). Given the structure of SEPAL, this results in zero-embeddings for entities not belonging to the core connected component at propagation time. Interestingly, the presence of zero-embeddings affects downstream scores very little, likely because most downstream entities belong to the core connected component and are thus not impacted by this.
  2. 2.
    METIS does not scale as well as BLOCS in terms of CPU memory. On our hardware, SEPAL + METIS could not scale to graphs larger than YAGO4.5 (32M entities), and therefore, BLOCS is indispensable for very large knowledge graphs.
E.4 Effect of BLOCS’ stopping diffusion threshold
Figure 13: Sensitivity analysis to parameter $h$. Effect of varying $h$ on BLOCS’ execution time, replication factor (average number of outer subgraphs containing a given entity), and maximum subgraph size. A lower bound of the domain of $h$ values to explore for a dataset is given by the proportion of nodes that are neighbors to super-spreaders. Indeed, as BLOCS’ first step is to assign super-spreaders neighbors, if $h$ is smaller than this value, BLOCS completely skips the diffusion phase. For YAGO4.5 and YAGO4, this results in sharp variations of the replication factor or the maximum subgraph size, respectively. The maximum size parameter $m$ was set to 400k for YAGO3, and 4M for others.
Figure 14: Outer subgraphs are well connected to the core. The left plot gives the percentage of outer entities that are directly connected to the core subgraph. The right plot gives the percentage of edges that come from the core subgraph among all the edges coming to a given outer subgraph, showing the amount of information transferred from the core to outer subgraphs relative to outer-outer communication.
In the BLOCS algorithm, the hyperparameter $h$ controls the moment of the switch from diffusion to dilation. For $h\in(0,1)$, BLOCS stops diffusion once the proportion of entities of the graph assigned to a subgraph is greater than or equal to $h$.
Figure 13 shows that increasing $h$ tends to increase the execution time and the overlap between subgraphs. Higher overlaps can be preferable to enable the information to travel between outer subgraphs during embedding propagation. However, a high overlap also incurs additional communication costs because the embeddings are moved several times from CPU to GPU, increasing SEPAL’s overall execution time.
E.5 How distant from the core are the outer entities?
Figure 14 shows that outer entities are in average very close to the core subgraph. This is due to the fact that the core contains the most central entities, and to the scale-free nature of knowledge graphs.
Appendix F Further analysis of SEPAL
F.1 Execution time breakdown
Here, we present the contribution of each part of the pipeline to the total execution time. Specifically, we break down our method into four parts:
  1. 1.
    core subgraph extraction;
  2. 2.
    outer subgraphs generation (BLOCS);
  3. 3.
    core embedding;
  4. 4.
    embedding propagation.
Figure 15: SEPAL’s execution time breakdown. Execution time of SEPAL’s different components, for the best-performing configurations of SEPAL on downstream tasks (see Table 9 for hyperparameters values).
Figure 15 shows the execution time of the different components of SEPAL. It includes six large-scale knowledge graphs, for which the execution times have the same order of magnitude. The results reveal that most of the execution time is due to the core embedding and embedding propagation phases, while the core extraction time is negligible.
Four key factors influence SEPAL’s execution time during the four main steps of the pipeline:
  1. 1.
    The core selection strategy: the degree-based selection is faster than the hybrid selection. For the hybrid selection, the factor that influences the speed the most is the number of distinct relations.
  2. 2.
    The diameter of the knowledge graph: graphs with large diameters call for more dilation steps during BLOCS’ subgraph generation, and dilation is more costly than diffusion because it requires checking node assignments. This explains why adding the taxonomies to YAGO4 and YAGO4.5 drastically reduces the time required to run BLOCS, as shown on Figure 15.
  3. 3.
    The core subgraph size: the more triples in the core subgraph, the longer the core embedding. This explains the wide disparities between the core embedding times on Figure 15, despite all the core subgraphs having roughly the same number of entities: YAGO4 core subgraph is more dense (33M triples), compared to YAGO4.5 (7M triples), for instance. The core embedding time also depends on hyperparameters such as the number of training epochs.
  4. 4.
    The total number $N$ of entities in the graph: this number determines the size of the embedding matrix. The communication cost of moving embedding matrices from CPU to GPU, and vice versa, accounts for most of the propagation time, and increases with $N$. It also increases with the amount of overlap between the outer subgraphs produced by BLOCS, explaining the differences in propagation time between YAGO4.5 and YAGO4.5 + taxonomy for instance.
The number of propagation steps $T$ has little impact on the embedding propagation time. The reason for this is that much of this time stems from the communication cost of loading the embeddings onto the GPU, and not from performing the propagation itself.
F.2 SEPAL’s hyperparameters
F.2.1 List of SEPAL’s hyperparameters
Here, we list the hyperparameters for SEPAL, and discuss how they can be set. Table 9 gives the values of those that depend on the dataset.
  • Proportion of core nodes $\eta_{n}$: the idea is to select it large enough to ensure good core embeddings, but not too large so that core embeddings fit in the GPU memory. Figure 16 shows the experimental effect of varying this parameter;
  • Proportion of core edges $\eta_{e}$: increasing it at the expense of $\eta_{n}$ (to keep the core size within GPU memory limits) improves the relational coverage, but reduces the density of the core. Sparser core subgraphs tend to deteriorate the quality of SEPAL’s embeddings for feature enrichment (Figure 20). However a good relational coverage is essential for better link-prediction performance;
  • Stopping diffusion threshold $h$: it depends on the graph structure, and tuning is done empirically by monitoring the proportion of unassigned entities during the BLOCS algorithm: $h$ is chosen equal to the proportion of assigned entities at which BLOCS starts to stagnate during its diffusion regime. In practice, Figure 13 and Figure 17 shows that as long as $h$ is chosen greater than the proportion of entities that are neighbors of super-spreaders, the algorithm is not too sensitive to this parameter (otherwise, it skips the diffusion phase, which can be detrimental);
  • Number of propagation steps $T$: it is chosen high enough to ensure reaching the remote entities (otherwise, they will have zeros as embeddings). Taking $T$ equal to the graph’s diameter guarantees that this condition is fulfilled. However, for graphs with long chains, this may slow down SEPAL too much. In practice, setting $T$ at 2–3 times the Mean Shortest Path Length (MSPL) usually embeds most entities effectively;
  • Propagation learning rate $\alpha$: it controls the proportion of self-information relatively to neighbor-information during propagation updates. For the DistMult model, this parameter has no effect during the first propagation step, when an entity is reached for the first time because outer embeddings are initialized with zeros and the neighbors’ message is normalized. In practice, the embedding of an outer entity can reach in one step a position very close to its fix point, and thus this parameter does not have much effect (Figure 18). For our experiments we typically use $\alpha=1$;
  • Subgraph maximum size $m$: the idea is to use the largest value for which it is possible to fit the subgraph’s embeddings in the GPU memory. We use $4\cdot 10^{4}$ for Mini YAGO3, $4\cdot 10^{5}$ for YAGO3, $2\cdot 10^{6}$ for WikiKG90Mv2, and $4\cdot 10^{6}$ for the other knowledge graphs;
  • Embedding dimension $d$: we use $d=100$ (except for complex embedings, where $d=50$ to keep the same number of parameters);
  • Number of epochs for core training $n_{\text{epoch}}$: see Table 9;
  • Batch size for core training $b$: see Table 9;
  • Optimizer for core training: we use the Adam optimizer with learning rate $\text{lr}=1\cdot 10^{-3}$;
  • Number $p$ of negative samples per positive for core training: we use $p=100$ (Table 9).
Table 9: Hyperparameter search space for SEPAL, and best values (in bold) for each knowledge graph on downstream tasks. The best values are those that gave the best average performance on the 4 validation tables and that were used to get the results in Figure 2 and Figure 3.
Dataset Parameter Grid (best in bold)
Mini YAGO3 $\eta_{n}$$\eta_{e}$ 0.05/—, 0.2/—, 0.025/0.005, 0.3/0.05
$h$ 0.6, 0.65, 0.7, 0.75, 0.8
$T$ 2, 5, 10, 15, 25
$n_{\text{epoch}}$ 12, 25, 50, 60, 75
$b$ 512
pp 1, 100, 1000
YAGO3 $\eta_{n}$$\eta_{e}$ 0.05/—, 0.1/—, 0.025/0.015, 0.3/0.05
$h$ 0.55, 0.6, 0.65, 0.7, 0.75, 0.77
$T$ 2, 5, 10, 15, 25
$n_{\text{epoch}}$ 16, 18, 25, 45, 50, 75
$b$ 2048, 4096, 8192
pp 1, 100, 1000
YAGO4.5 $\eta_{n}$$\eta_{e}$ 0.03/—, 0.015/0.01, 0.05/0.03
$h$ 0.4, 0.5, 0.6
$T$ 2, 5, 10, 15, 25, 50
$n_{\text{epoch}}$ 16, 24, 75
$b$ 8192, 16384
pp 1, 100
YAGO4.5+T $\eta_{n}$$\eta_{e}$ 0.03/—, 0.015/0.005, 0.04/0.015
$h$ 0.8
$T$ 2, 5, 10, 15, 20, 25
$n_{\text{epoch}}$ 32, 75
$b$ 8192
pp 100
YAGO4 $\eta_{n}$$\eta_{e}$ 0.03/—, 0.015/0.005, 0.012/0.025
$h$ 0.55
$T$ 2, 5, 10, 15, 20, 25
$n_{\text{epoch}}$ 4, 28, 75
$b$ 8192, 65536
pp 1, 100
YAGO4+T $\eta_{n}$$\eta_{e}$ 0.02/—, 0.01/0.005
$h$ 0.45, 0.8
$T$ 10, 20
$n_{\text{epoch}}$ 32, 75
$b$ 8192, 65536
pp 1, 100
Freebase $\eta_{n}$$\eta_{e}$ 0.02/—, 0.01/0.005
$h$ 0.55
$T$ 15
$n_{\text{epoch}}$ 24
$b$ 8192
pp 100
WikiKG90Mv2 $\eta_{n}$$\eta_{e}$ 0.02/—
$h$ 0.92
$T$ 10
$n_{\text{epoch}}$ 12
$b$ 8192
pp 100
Figure 16: Effect of core proportion $\eta_{n}$ on SEPAL’s performance, with the degree-based core selection strategy.
Figure 17: Effect of stopping diffusion threshold $h$ on SEPAL’s performance.
Figure 18: Effect of propagation learning rate $\alpha$ on SEPAL’s downstream performance.
Figure 19: Effect of number $T$ of propagation steps on SEPAL’s downstream performance.
F.2.2 Experimental study of hyperparameter effect
Figure 16, Figure 17, Figure 18 and Figure 19 investigate the sensitivity of SEPAL to different hyperparameters. The hyperparameter that seems to impact the most SEPAL’s performance is the core proportion $\eta_{n}$. Indeed, Figure 16 shows that increasing $\eta_{n}$ tends to improve embedding quality for downstream tasks. However, the effect seems to be plateauing relatively fast for YAGO3 (not much improvement between $\eta_{n}=5\%$ and $\eta_{n}=10\%$). For other datasets (YAGO4.5, YAGO4), it is not possible to explore larger values of $\eta_{n}$ because the core subgraph would not fit in the GPU memory. Moreover, decreasing $\eta_{n}$ makes SEPAL run faster, as the core embedding phase accounts for a substantial share of the total execution time (Figure 15). There is, therefore, a trade-off between time and performance.
Core selection strategy
Table 10: Effect of core selection strategies. Number of entities and triples inside the core subgraph and proportion of the full graph they represent (in parentheses). $\eta_{n}$ and $\eta_{e}$ are the hyperparameters for nodes and edges, respectively. Column #Rel gives the number of relation types present in the core compared to the total number of relation types in the knowledge graph. We highlight in red the cases where some relations are missing. Column Time gives the measured computation time for core selection.
Strategy $\boldsymbol{\eta_{n}}$ $\boldsymbol{\eta_{e}}$ #Rel #Entities #Triples Time
Degree 5% - 37/37 126k (4.9%) 1.0M (18.5%) 17s
YAGO3 Hybrid 2.5% 1.5% 37/37 121k (4.7%) 733k (13.1%) 20s
Degree 3% - 62/62 932k (2.9%) 7.2M (9.6%) 2min
YAGO4.5 Hybrid 1.5% 1% 62/62 1.1M (3.3%) 5.7M (7.5%) 6min
Degree 3% - 61/76 1.1M (3.0%) 33M (13.4%) 8min
YAGO4 Hybrid 1.5% 0.5% 76/76 1.4M (3.8%) 28M (11.1%) 11min
Degree 3% - 64/64 1.5M (3.0%) 13M (9.9%) 4min
YAGO4.5+T Hybrid 1.5% 0.5% 64/64 1.2M (2.5%) 8.3M (6.5%) 5min
Degree 2% - 64/78 1.3M (2.0%) 41M (12.8%) 9min
YAGO4+T Hybrid 1% 0.5% 78/78 1.5M (2.3%) 32M (10.1%) 12min
Degree 2% - 5,363/14,665 1.7M (2.0%) 15M (4.4%) 9min
Freebase Hybrid 1% 0.5% 14,665/14,665 1.9M (2.3%) 14M (4.1%) 4h
Degree 2% - 886/1,387 1.8M (2.0%) 62M (10.4%) 17min
WikiKG90Mv2 Hybrid 0.4% 0.7% 1,387/1,387 2.0M (2.2%) 37M (6.2%) 48min
Degree-based selection: The simple degree-based core selection strategy is convenient for two reasons:
  1. 1.
    Degree is inexpensive to compute, ensuring the core extraction phase to be fast (see Figure 15);
  2. 2.
    It yields very dense core subgraphs. Indeed, while they contain $\eta_{n}$% of the entities of the full graph, they gather around $4\eta_{n}$% of all the triples (Table 10). This allows the training on the core to process a substantial portion of the knowledge-graph triples, resulting in richer representations.
Hybrid selection: However, a problem with the degree-based selection is that some relation types may not be included in the core subgraph. To deal with this issue, SEPAL also proposes a more complex hybrid method for selecting the core subgraph, that incorporates the relations. It proceeds in four main steps:
  1. 1.
    Degree selection: Sample the nodes with the top $\eta_{n}$ degrees.
  2. 2.
    Relation selection: Sample the edges with the top $\eta_{e}$ degrees (sum of degrees of head and tail) for each relation type, and keep the corresponding entities.
  3. 3.
    Merge: Take the union of these two sets of entities.
  4. 4.
    Reconnect: If the induced subgraph has several connected components, add entities to make it connected. This is done using a breadth-first search (BFS) with early stopping from the node with the highest degree of each given connected component (except the largest) to the largest connected component. For each connected component (except the largest), a path linking it to the largest connected component is added to the core subgraph.
This way, each relation type is guaranteed to belong to the core subgraph, by design. Table 10 confirms this experimentally, even for Freebase and its 14,665 relation types. This method features two hyperparameters $\eta_{n}$ and $\eta_{e}$, the proportions for node and edge selections, controlling the size of its output subgraph. The values we used are provided in Table 10 for each dataset.
Regarding performance, Figure 20 shows that the hybrid strategy slightly underperforms the degree-based approach on four real-world downstream tasks. Indeed, the hybrid approach enhances relational coverage but, as a counterpart, yields a sparser core subgraph (see #Triples in Table 10), which is detrimental to downstream performance. However, one can adjust the values of $\eta_{n}$ and $\eta_{e}$ to control the trade-off between downstream performance and relation coverage, depending on the use case. For instance, for the link prediction task, a better relational coverage (greater $\eta_{e}$) improves the performance (Section˜C.6).
Figure 20: Performance of SEPAL+DistMult for the two different core selection strategies. We use the hyperparameters of Table 10. The simpler degree-based selection strategy runs faster and performs better on downstream tasks.
Table 11: Effect of core selection strategy: PageRank vs Degree. We compare the performance of SEPAL+DistMult on real-world downstream tasks when using either degree or PageRank as a centrality measure for core selection. We report the average normalized R2 score across the four downstream tasks, along with the total execution time (in parentheses).
Dataset Core defined by Degree Core defined by PageRank
YAGO3 0.783 (11m 20s) 0.742 (9m 23s)
YAGO4 0.817 (6h 20m) 0.861 (4h 53m)
YAGO4+T 0.815 (10h 9m) 0.881 (8h 16m)
YAGO4.5 0.949 (4h 11m) 0.925 (4h 3m)
YAGO4.5+T 0.923 (2h 47m) 0.912 (2h 40m)
Freebase 0.917 (5h 58m) 0.919 (5h 58m)
WikiKG90Mv2 0.898 (20h 31m) 0.902 (23h 59m)
Exploring other centrality measures
Here, we compare degree with PageRank as a centrality measure for the core selection (keeping all other hyperparameters unchanged). Like the degree, the PageRank can be computed very efficiently on huge graphs using sparse matrix multiplications.
Table 11 reports results on the real-world downstream tasks. Regarding performance, it seems that the difference between degree and PageRank depends on the graph’s structure. Specifically, we observe that for YAGO4, YAGO4+T, Freebase, and WikiKG90Mv2, PageRank improves the performance. This echoes the results of Figure 14, which show that in these four specific KGs, contrary to the others, the degree-based core selection yields cores that are connected through fewer edges to some outer subgraphs. We can therefore conjecture that for these KGs, PageRank improves information flow and mitigates issues like oversquashing during propagation, ultimately increasing performance. Regarding computational cost, we see that SEPAL with PageRank usually runs slightly faster than with degree. This is because PageRank yields slightly sparser cores, leading to faster core training.
F.3 Ablation study: SEPAL without BLOCS
Here, we study the effect of removing BLOCS from our proposed method. On smaller knowledge graphs, SEPAL can be used with a simple core subgraph extraction and embedding followed by the embedding propagation. This ablation reveals the impact of BLOCS on the model’s performance. Figure 21 shows that adding BLOCS to the pipeline on graphs that would not need it (because they are small enough for all the embeddings to fit in GPU memory) does not alter performance. Additionally, BLOCS brings scalability. By tuning the maximum subgraph size $m$ hyperparameter, one can move the blue points horizontally on Figure 21 and choose a value within the GPU memory constraints. There is a trade-off between decreasing GPU RAM usage (i.e., moving the blue points to the left) and increasing execution time, as fewer entities are processed at the same time.
Figure 21: Ablation study: BLOCS scales SEPAL memory-wise. Normalized R2 scores aggregated across evaluation datasets (movie revenues, US accidents, US elections, housing prices) for SEPAL with and without BLOCS are plotted against GPU RAM usage. BLOCS preserves performance for a given knowledge graph while drastically reducing memory pressure on GPU RAM. Without BLOCS, the GPU runs out of memory for YAGO4 and YAGO4.5.
Figure 22: Comparison between SEPAL and its base embedding model. a) Computation time per training iteration. For a given training configuration, SEPAL is up to 30 times faster than its base embedding algorithm DistMult. b) Learning curves. SEPAL achieves strong downstream performance much quicker than DistMult. For both plots, we only vary the number of epochs and fix the following parameters for DistMult’s training and SEPAL’s core training: $p=1$, $\text{lr}=1\cdot 10^{-3}$, $b=512$ for Mini YAGO3 and $b=2048$ for YAGO3. For SEPAL, we use the degree-based core selection with $\eta_{n}=5\%$.
F.4 Speedup over base embedding model
Figure 22 demonstrates that SEPAL can accelerate its base embedding model by more than a factor of 20, while also boosting its performance on downstream tasks.
Moreover, the speedup increases with the number of training epochs, as SEPAL’s constant-cost steps (core extraction, BLOCS, and propagation) are amortized when core training gets longer. Indeed, each additional epoch is cheaper with SEPAL than with DistMult, since training is done only on the smaller core rather than the full graph.
Appendix G Discussion
G.1 Comparison to prior work
G.1.1 Comparison to DistMult-ERAvg
albooyeh2020out propose an aggregation that is similar to SEPAL’s aggregation during the propagation phase, however the two methods optimize the embeddings in two very distinct ways:
  • albooyeh2020out introduce propagation within the standard link prediction pipeline: during training, for each triple $(v,r,u)$, they occasionally (with probability $p$) replace one entity’s embedding with an aggregated version (e.g., $\theta_{u}\cdot\theta_{r}$) before computing the plausibility score. However, training still relies on negative sampling and a classic link prediction loss, and thus optimizes for local triple-level contrasts like traditional KGE methods.
  • SEPAL, in contrast, explicitly separates the optimization objective: embeddings for a small core are trained with a classic KGE objective, and then relation-aware propagation is used across the rest of the graph, without negative sampling. This distinction is crucial: SEPAL removes the need for negative sampling on typically 95 to 99% of the graph, enabling both improved scalability and alignment properties beneficial for downstream tasks.
Moreover, DistMult-ERAvg is not designed for very large graphs. On the datasets considered in this paper, it fails with out-of-memory errors on all the graphs except Mini YAGO3 (129k entities, 1.1M triples).
Table 12: Comparison between SEPAL and DistMult-ERAvg on Mini YAGO3. We report the normalized mean cross-validation score (R2) across the four real-world downstream tasks, along with the total training time. SEPAL outperforms DistMult-ERAvg while being significantly faster.
Method Housing prices Movie revenues US accidents US elections Time
DistMult-ERAvg 0.149 0.124 0.444 0.916 2h 41m
SEPAL 0.276 0.159 0.548 0.929 5m
Table 12 reports performance and runtime on Mini YAGO3, showing that SEPAL is 32 times faster than DistMult-ERAvg while achieving consistently better scores. This is expected since DistMult-ERAvg follows the classic optimization loop with negative sampling, gradient computations, and parameter updates, and therefore inherits the limitations of traditional KGE methods. The strength of DistMult-ERAvg lies in out-of-sample embedding computation, which SEPAL also supports (see Section˜G.3), but with greater scalability.
G.1.2 Comparison to NodePiece
SEPAL shares with NodePiece the fact that it embeds a subset of entities. Parallels can be drawn between: a) the anchors of NodePiece and the core entities of SEPAL; b) the encoder function of NodePiece and the embedding propagation of SEPAL. Yet, our approach differs from NodePiece in several ways.
Neighborhood context handling
Both methods handle completely differently the neighborhood of entities. NodePiece tokenizes each node into a sequence of $k$ anchors and $m$ relation types, where $k$ and $m$ are fixed hyperparameters shared by all nodes. If the node degree is greater than $m$, NodePiece downsamples randomly the relation tokens, and if it is lower than $m$, [PAD] tokens are appended; both seem sub-optimal. In contrast, SEPAL accommodates any node degree and uses all the neighborhood information, thanks to the message-passing approach that handles the neighborhood context naturally.
Additionally, NodePiece’s tokenization relies on an expensive BFS anchor search, unsuitable for huge graphs. On our hardware, we could not run the vanilla NodePiece (PyKEEN implementation) on graphs bigger than Mini YAGO3 (129k entities). For YAGO3 and YAGO4.5, we had to run an ablated version where nodes are tokenized only from their relational context (i.e., $k=0$, studied in the NodePiece paper with good results), to skip the anchor search step.
Training procedure
At train time, NodePiece goes through the full set of triples at each epoch to optimize both the anchors’ embeddings and the encoder function parameters, necessitating many gradient computes and resulting in long training times for large graphs. On the contrary, SEPAL performs mini-batch gradient descent only on the triples of the core subgraph, which provides significant time savings. To illustrate this, Figure 23 compares the performance of SEPAL and vanilla NodePiece on Mini YAGO3, showing that SEPAL outperforms NodePiece on downstream tasks while being nearly two times quicker.
Figure 23: Comparing SEPAL with NodePiece on Mini YAGO3.
Embedding propagation to non-anchor/non-core entities
To propagate to non-anchor entities, NodePiece uses an encoder function (MLP or Transformer) that has no prior knowledge of the relational structure of the embedding space, and has to learn it through gradient descent. On the contrary, SEPAL leverages the model-specific relational structure to compute the outer embeddings with no further training needed.
G.2 Communication costs of SEPAL
SEPAL optimizes data movement
In modern computing architectures, memory transfer costs are high, amounting to much computation, and the key to achieve high operation efficiency is to reduce data movement [mutlu2022modern].
For distributed methods such as PBG or DGL-KE, parallelization incurs additional communication costs due to two factors: 1) Learned parameters shared across workers (relation embeddings) require frequent synchronization 2) Entities occurring in several triples belonging to different buckets have their embeddings moved several times from CPU to GPU, and this for every epoch. This latter effect can be mitigated by better partitioning, but remains significant.
SEPAL avoids much of the communication costs by keeping the core embeddings on GPU memory throughout the process. These core embeddings are the ones that would move the most in distributed settings, because they correspond to high-degree entities involved in many triples. Moreover, SEPAL’s propagation loads each outer subgraph only once on the GPU, contrary to other methods that perform several epochs. This significantly reduces data movement for outer embeddings: empirically, they only cross the CPU/GPU boundary twice on average (see Table 13).
Table 13: Empirical average number of memory transfers between CPU and GPU for outer embeddings across datasets.
Dataset $\boldsymbol{x_{\mathrm{outer}}}$
Mini YAGO3 1.20
YAGO3 3.07
YAGO4.5 7.47
YAGO4.5+T 1.94
YAGO4 2.29
YAGO4+T 1.27
Freebase 2.09
WikiKG90Mv2 1.84
Average 2.65
Estimating I/O
Communication costs occur in SEPAL during the propagation phase, where the embeddings of each of the subgraphs generated by BLOCS have to be loaded on the GPU, subgraph after subgraph. We analyze the number $x$ of back-and-forth of a given embedding between CPU and GPU memory. The optimal value is $x=1$, meaning the embeddings are transferred only once. A detailed breakdown follows:
  • For core embeddings: core embeddings remain at all times on the GPU memory. They are only moved to the CPU once at the end, to be saved on disk with the rest of the embeddings. Therefore, the data movement of core embeddings is optimal: $x_{\mathrm{core}}=1$.
  • For outer embeddings: SEPAL loads each outer subgraph only once to the GPU. Consequently, the average number $x_{\mathrm{outer}}$ of memory transfers for outer embeddings corresponds to the average number of subgraphs in which a given outer entity appears. Table 13 reports empirical values of $x_{\mathrm{outer}}$ across datasets, ranging from 1.20 to 7.47, with an overall average of 2.65. This redundancy arises from subgraph overlap, which is directly influenced by the $h$ hyperparameter in BLOCS: smaller values of $h$ yield fewer diffusion steps (and more dilation steps), leading to reduced subgraph overlap and, hence, fewer memory transfers per embedding.
The optimal value of $x=1$ can only be achieved in the case where the entire graph fits in GPU memory. Every approach that scales beyond GPU RAM limits has its communication overheads, i.e., $x>1$.
For comparison, distributed training schemes that load buckets of triples to GPU iteratively exhibit significantly higher communication costs. In such settings, the number of memory transfers for the embedding of an entity $u$ is given by $x=n_{\mathrm{buckets}}(u)\times n_{\mathrm{epochs}}$ where $n_{\mathrm{buckets}}(u)$ is the number of buckets that contain a triple featuring entity $u$, and $n_{\mathrm{epochs}}$ is the number of epochs. That is at least one or two orders of magnitude greater than what SEPAL achieves in terms of data movement. Indeed, knowledge-graph embedding methods are usually trained for a few tens if not a few hundreds of epochs, so $x$ is much bigger than 10, not even taking into account $n_{\mathrm{buckets}}(u)$ that can be large, depending on the graph structure and on the quality of the graph partitioning.
G.3 Outlook on continual learning
Table 14: Statistics of the two versions of YAGO3 used to illustrate SEPAL’s suitability for continual learning.
Version #Entities #Relations #Triples
YAGO3-2014 2,570,716 37 5,585,004
YAGO3-2022 4,546,966 37 14,691,781
The modular nature of SEPAL makes it well-suited for continual learning scenarios, where new entities are added to the knowledge graph over time. Indeed, new embeddings can be computed without retraining from scratch, via a few additional propagation steps, as long as the relations remain unchanged. To demonstrate this, we use two versions of the YAGO3 knowledge graph: the original 2014 release, and the 2022 revived version. Table 14 gives the statistics of these two datasets.
Table 15: Continual learning experiment on YAGO3. Average normalized R2 scores across the four real-world downstream regression tasks (movie revenues, US accidents, US elections, housing prices) and total runtime for different methods. SEPAL-CL denotes SEPAL in the continual learning setting, where new entities are embedded via propagation.
YAGO3 version Method Average performance Runtime
2014 SEPAL 0.836 0h 11m 20s
2022 DistMult 0.884 11h 05m 36s
2022 SEPAL 0.988 1h 05m 07s
2022 SEPAL-CL 0.962 0h 01m 08s
We adapt SEPAL to this continual learning setting by: (1) initializing the embeddings of YAGO3-2022 using embeddings precomputed on YAGO3-2014, (2) propagating embeddings for 5 additional steps to update existing nodes and embed new entities. We denote this method SEPAL-CL in the results.
Table 15 shows that:
  • For downstream applications, YAGO3-2022 brings value compared to YAGO3-2014. Indeed, for all the methods considered, the embeddings learned on the 2022 dataset score higher than the strongest method on YAGO3-2014, SEPAL.
  • SEPAL trained from scratch is 10 $\times$ faster than DistMult on YAGO3-2022.
  • SEPAL-CL is 57 $\times$ faster than SEPAL trained from scratch on YAGO3-2022, and 587 $\times$ faster than DistMult.
  • SEPAL-CL clearly outperforms DistMult, and almost matches the performance of SEPAL trained from scratch, even in this very challenging scenario (8 years between the two versions of the graph), where the size of the graph has doubled in terms of entities, and tripled in terms of triples. In a real-life application, we could imagine embeddings recomputed monthly at very low cost.
G.4 Broader impacts
SEPAL may reflect the biases present in the training data. For instance, Wikipedia, from which YAGO is derived, under-represents women [reagle2011gender]. We did not evaluate how much our method captures such biases. We note that the abstract nature of embeddings may make the biases less apparent to the user; however, this problem is related to embeddings and not specific to our method. It may be addressed by debiasing techniques [bolukbasi2016man, fisher2020debiasing] for which SEPAL could be adapted.